A Complete Bibliography of Econometrics and Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

24 September 2024 Version 1.18

Title word cross-reference

5^{th} [AKW20]. A [CCCW23]. F [Cha23]. g [GP23, HPW23]. I [Ber20]. k [Fer24, LEP18]. l_{0} [ZY24]. L_{4} [Dem21]. M [OSSRK19]. Z^{2} [DF19]. O(n) [Dem21]. t [MBM17, PCP20].

-class [LEP18]. -Means [Fer24]. -optimal [CCCW23]. -priors [GP23, Ber20]. -quantile [OSSRK19].

1-130 [Ano17a]. 1-136 [Ano20e]. 1-148 [Ano17b]. 1-158 [Ano19e, Ano20f]. 1-164 [Ano18a]. 1-168 [Ano17c, Ano20g]. 1-170 [Ano19f, Ano19g]. 1-172 [Ano21h]. 1-188 [Ano18b]. 1-196 [Ano20h]. 1-200 [Ano17d]. 1-216 [Ano19h]. 1-250 [Ano18c]. 19 [BCCI24]. 1st [CK21b].

2017 [Ano17b, Ano17d, Ano17c, Ano17a]. 2018 [Ano18b, Ano18a, Ano18c]. 2019 [Ano19f, Ano19g, Ano19e, Ano19h]. 2020 [Ano20f, Ano20h, Ano20e, Ano20g]. 2021 [Ano21e, Ano21h, Ano21f, Ano21g]. 2022 [Ano22e, Ano22g, Ano22f]. 2023 [Ano23f, Ano23e, Ano23h, Ano23g]. 2024 [Ano24f, Ano24h, Ano24e, Ano24g]. 20th [CK23]. 2nd [CK21c, FKKP22, IL22]. 2SLS [LEP18].

50 [CG19]. 500 [HO18].

accepting [DF19]. Accurate [Ron20]. Achieving [FY19]. across [Bla22]. Adaptive [BS21, OSSRK19, WW22, FR17, GF19, YW23]. AdaVol [WW22]. Additive [ZY24, AMK22, Yos18]. adjusted [BSTS22]. Advances [CK21b, CK21c]. affine [HS20]. Against [CY21, KSS23, Mar17, MGF18]. age [Beu23]. Aggregation [dBCR21, Cha22]. agricultural [Czu19]. algorithm [AFKL21, BB21, Dem21, Flo18, For17]. algorithms [CK24, FGV21]. allocation [GP17]. allows [TB20]. Alternative [Hay19, HK22, MGF18]. among [HVBJ17]. analyses [BHKX17]. Analysis [FKKP22, KR22, SSL23, ZPY^{+}24, CKL^{+}23, CS22, DJF20, GP23, GJ19, Hay24, KOPS17, LS21, MTV17, PS19a, RC22, AGE17]. analytic [ÖA17]. analytical [NW19]. Analytically [CCCW23]. analyzers [MBM17]. Analyzing [KLMO22]. annals [AKW20]. annealing [Flo18]. anniversary [CK23]. Annual [BD22, BHKX17]. annuity [ZMG18]. anti [AMS23]. anti-mode [AMS23]. antipersistent [ADG24]. Application [Art24, Asa23, RPRH19, BFS18, CWL23, CK22, DJF20, FZF19, GM18, HO18, KKW17, KJ20, Sha17, SAK21, ZZ19]. Applications [VQKS24, AL17, Fer24, FFO20, Gui18]. applied [De 20, Lux22, Mac21b]. Approach [ZYL22, BP22, CZJ^{+}17, Czu19, DFS20, FH19, GD22, HT23, Kar18, KO17, PS19b, SFdB17, VS17]. Approaches [KSS23]. Approximating [BKP18]. Approximation [LSW24, Dem21, FKM24]. approximations [dZBMRV24, LEP18]. April [Ano17b, Ano19f, Ano20f, Ano21e, Ano23f, Ano24f]. Arbitrage [PS23, Mac21a]. arbitrage-free [Mac21a]. Arbitrary [RAC^{+}24]. ARCH [CA17, SSL23]. ARCH-M [SSL23]. area [Mor24]. arising [QD19]. ARMA [DF19, KKW17, LDL21, SWS22]. ARMA-type [DF19]. ARTFIMA [VQKS24]. Assessing [BS18]. associated [Beu23]. assumption [HP21]. Asymmetric [ACG21, MCRV20, Pao17, FH21]. Asymptotic [BCR24, Mél22, ST24b]. Asymptotics [FS22]. attitude [STI20]. augmented [MPG18, WFSJ23]. auto [De 20]. auto-correlation [De 20]. Automatic [KR22]. autoregressions [FY19, LN17]. Autoregressive [SM20, CA17, HKTZ19, HV21, MTV17, SZH19]. average [NW19, PS19b]. Averaging [HPW23, PL24]. Avoiding [RAC^{+}24].

Babai [CCW24]. band [BS18]. bands [GFED22, LSBW20]. bandwidth [AO17, WL21]. Based [GM22, MW20, AADN18, CCW24, CR20, CKP21, CS22, ELS23, FHT20, MCRV20, MGF18, MMS^{+}24, dVC18]. Bayes [HPW23]. Bayesian [AMS23, CLNP23, CWL23, CW23, COSW17, CNBN22, Czu19, DJF20, FHT20, FY19, GP23, GW17, HGFO20, KC21, LCD21, LS21, MTV17, PS19a, PCP20, SG17, SSL23, THS18, VS17, WFSJ23, ZYL22, ZDY22]. behaviour [PPDZ18]. Benjamini [Döh18]. best [Dem21]. Beta [Cha23]. Beta-statistics [Cha23]. between [Czu19]. beyond [CK24]. Bias [BKH22, DGK22, FKM24, FH21, BSTS22, GEH19, LEP18]. bias-adjusted [BSTS22]. Bias-corrected [BKH22]. bias-reducing [GEH19]. Big [Mar17]. Binary [AADN18, CF21, COySS23, FM17, BBP17, PB22]. binomial [ELS23]. Biostatistics [GR23]. Bitcoin [PCP20]. bivariate [PCP20]. block [KO20]. blocked [GG23]. Blockwise [MOCB21]. Board [Ano19a, Ano19d, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano24a, Ano24b, Ano24c, Ano24d]. book [PPDZ18]. boosting [AMK22]. Bootstrap [Ang20, ZP21, BS19, GRC17, Mac23]. Bootstrapping [Art24]. both [NABM23]. boundary [Kir20]. brain [HGFO20]. Break [BS21]. breaks [AÖ21, MTV17]. Brief [AM23]. Brownian [LS20]. bubbles [HV21]. Business [GOS24]. BVARs [FHT20].

calibration [SMSC21]. Carlo [BLvV21, KC21, LCD21, WCD22]. case [DKA17]. categorical [HVBJ17]. causal [HV21, LEB17, SM19]. causal-noncausal [HV21]. causality [AS19, BS18]. censored [SV18]. censoring [NABM23]. Central [HKTZ19]. CFEnetwork [AKW20]. chain [ELS23, FS22]. Change [BHKX17, ELS23, GHK18, CK24]. changepoints [Mac21b]. changes [JM21, SAK21]. characterizations [Ley23]. checking [LDL21]. childhood [CKL^{+}23]. Chinese [THS18]. choice [AADN18]. choice-based [AADN18]. Cholesky [MMS^{+}24, SOLP17]. Cholesky-based [MMS^{+}24]. Choosing [ACG21]. circular [AMS23]. class [FGV21, For17, LEP18, PV17, QD19, Ric20]. Classification [PL24]. Classifiers [AM23, ANL^{+}21]. cleaning [MPG18]. Closed [PS19b]. Closed-form [PS19b]. closer [TB20]. cluster [Mac23]. clustered [BBP17, BSTS22]. Clustering [KR22, FGV21, HVBJ17, SC22, ST24a]. CMStatistics [CK23]. coefficient [Bla22, CK21a, Czu19, SZH19]. cointegrating [AÖ21, ÖA17]. cointegration [VS21]. collinear [PS19a]. Combined [CA17, HLU20]. Combining [BCR24, Bay18, Cha23]. Commodity [KLMO22]. common [CRT19, CG19, DF21]. Comovements [GOS24]. Comparison [Lux22, CJG24, Egu18]. competitive [KO17]. component [BBS17]. components [ACG21, CG19, Max20, PRV17]. Composite [GM18, WCW18]. computation [BK18]. Computational [CK23, AKW20]. concentration [Sha17]. Concurrent [GM22]. condition [Mor24]. Conditional [LSY24, PB22, SM20, BB21, BP22, GD22, GSUC22, GGHQ23, Mor19]. conditionally [LDL21]. conditions [KPY18, RSV20]. confidence [GFED22, LSBW20, SFdB17, SMSC21]. confined [Kiv23]. Connectedness [GOS24]. connectivity [HGFO20]. Consistency [ST24b, GRC17]. consistent [ySL21]. constituent [HO18]. constrained [KT19]. constraints [WS18]. Constructing [CKL^{+}23, LSBW20]. consumer [SFdB17]. Consumption [BD22, RC22]. contaminated [FZF19]. context [Mar17]. Continuous [CF21, KALP24, AGMR18, KLS18]. continuous-time [KLS18]. controls [BK18]. Convergence [Kaw22, HSS18]. convex [GFED22]. Copula [KR22, KH19, ACL19, BM22, ELS23, GD22, GKD19, KC21, KJ20, SC22, SV18, ZDY22]. copula-based [ELS23]. copula-graphic [SV18]. Copulas [Smi23, FFO20, QD19, SS23]. corporate [GM18]. Corrected [LBWM22, BKH22]. Correcting [DGK22, De 20]. correction [Ang20, FKM24, FH21]. correlation [De 20, ST24b]. Correlations [BCR24]. Count [Fok24, LSY24, BFS18]. Covariance [Hir23, Li24, BBS17, Hay24, MOCB21, Mor19, SAK21]. Covariate [XY22]. Covariate-Varying [XY22]. Covariates [Bin18, For17, PB22]. COVID [BCCI24]. COVID-19 [BCCI24]. credit [DCVB22, JM21]. Criteria [RAC^{+}24]. criterion [KH19, Lam18]. cross [AÖ21, JM21, RC22]. cross-sectional [AÖ21, JM21, RC22]. cryptocurrencies [PCP20]. cure [DCVB22]. curves [BHKX17, GFED22, KPTV21, THS18]. CUSUM [WL21]. Cycle [GOS24]. Cyclical [VS21]. cylindrical [AL17].

d [BM22]. d-vines [BM22]. DAGs [BB21]. Daily [TWO24]. Data [CK24, COySS23, FKKP22, LSY24, PZK22, ZZZ20, ZPY^{+}24, AL17, ACL19, ASJLZ17, ADN18, BFS18, BBP17, BSTS22, BCJC17, BKH22, CdL20, Cha23, CNBN22, CKL^{+}23, CS22, DJF20, DFS20, DF21, FZF19, GOP18, HVBJ17, Hay19, HO18, HLU20, KPY18, Kiv20, KKW17, KOPS17, KJ20, LS20, Mac21b, MPG18, PB22, RC22, SC22, SAK21, Tan24, WCW18, ZZ19, Mar17]. data-cleaning [MPG18]. decision [YW23]. decrement [ZMG18]. defect [ANL^{+}21]. delay [BS18]. demand [BFS18, Ric20]. densities [FH19]. Density [ADN18, PZK22, BS19, CR20, Kaw22]. Dependence [OP24, AÖ21, CIO19, FFO20, GKO18, HO18, JM21, Kir20, KJ20, RC22, ZDY22]. Dependent [PS24, BI19, De 20, Egu18, ÖA17]. Designating [PPDZ18]. Designs [CCCW23, BSM23, GG23, YW23]. detect [FMU20]. Detecting [SAK21]. Detection [RAC^{+}24, RPRH19, ANL^{+}21, FS22, GHK18, GF19]. development [Hay24]. diagnostic [LDL21]. difference [KV24, LDL21, MGF18]. differences [Bla22, TB20]. Different [Lux22, LN17]. differential [HGFO20, MZ22]. Differentially [KALP24]. Differentiation [KR22]. diffusion [CHM24]. Dimension [BB24, MCNO17]. Dimensional [LBWM22, LDA23, ZY24, BBS17, CK22, FR17, HO18, LY19b, Mor19, ZZZ20]. Direct [MW20]. discontinuity [FH19]. discount [PS23]. Discovery [CDH23, Yos18]. discrete [Dem21, Döh18, STW18]. Discrimination [STW18]. dispersion [TB17]. distance [PV17]. distortion [ES18]. distress [FMU20]. distribution [AFKL21, Jia17, Ley23, Pao17]. Distributional [KSS23]. distributions [AMS23, BKP18, ES18, QD19]. disturbances [LEP18]. Divisible [ZHO24]. domain [FFO20]. Double [ACG21]. Driven [BLvV21, FM17]. DSGE [Ang20, FHT20, Mut18]. DSGE-based [FHT20]. Dynamic [HW24, HL23, Li24, LDA23, ZYL22, ACL19, BBS17, BKH22, GW17, HT23, Hu21, Kiv20, KJ20, KO20, Mor19, NW19, Sha17, SSL23, ZZ19]. Dynamics [Czu19, ZZ19].

EC [PBMS20]. Econometrics [CK21b, CK21c, AKW20, BHL^{+}17, COSW17, DHW20, CRG23, KVC17]. Economic [OFG22, HSS18]. economics [Hay24]. EcoSta [CK21b, CK21c, DHW20]. edge [DF19]. edge-effect [DF19]. Edgeworth [JLM24]. Editorial [Ano19a, CK23, GR23, Ano19d, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano24a, Ano24b, Ano24c, Ano24d]. effect [CS22, DF19, Mul18, TB20]. effective [Beu23]. effects [AGMR18, BBP17, BH24, CJG24, May22, NW19, Pol18, WFSJ23, ZZ19]. Efficient [ANL^{+}21, LCD21, SG17]. eigenvalues [BV24]. Embedded [XY22]. Empirical [Asa23, HPW23, BP22, BCJC17, Lam18]. endogeneity [KP17, Kiv23, KO20]. Endogenous [CF21, BP22]. Energy [RC22]. England [HKTZ19]. Ensembling [LYBH21]. environment [CZJ^{+}17]. equation [LS21]. equations [BP22, PBMS20]. equicorrelations [KO20]. Equity [ABC24]. equivalence [BB24]. Equivalent [HP21]. ergodic [CHM24]. error [De 20]. errors [NW19]. estimate [Kar18]. estimates [CR20]. Estimating [GQ19, KO17, THS18, BP22, CCW24, HK22]. Estimation [ASJLZ17, Asa23, CF21, CHM24, HSS18, Hir23, LBWM22, Li24, SWS22, SZH19, ZZ19, ZY24, AMS23, ADN18, AO17, CWL23, CMR18, DJF20, DFS20, De 20, ELS23, FHT20, Flo18, FH19, FH21, GEH19, GSUC22, Hay19, HS20, HLU20, KV24, KH19, LCD21, MCRV20, MPG18, Mor19, PRV17, PS23, PS19b, SM19, WS18]. estimator [ADGG20, BSTS22, CRT19, GGHQ23, SV18, SVG19, dVC18]. estimators [Art24, BKH22, ES18, Kaw22, LEP18, Mél22, ST24b]. Euclidean [MOCB21]. euro [Mor24, CG19]. Evaluating [DF21]. event [STW18]. Evidence [BLvV21]. Evolutionary [HVBJ17]. Exact [ZHO24]. exogeneity [May22]. expansions [JLM24]. expectation [BB21]. expectations [Ang20]. expected [BKP18, GGHQ23, HT23]. expectile [BHKX17]. expectiles [GSUC22]. explained [CJG24]. explanatory [KP17, TB20]. extended [CCW24]. extracted [GOP18]. extraction [AGE17]. extreme [ADGG20, ES18, GSUC22, RSV20]. extremes [Gui18].

Factor [GT23, KLMO22, LDA23, WFSJ23, ZYL22, DF21, HT23, KC21, MBM17, NW19, PS23]. Factor-augmented [WFSJ23]. Factors [ZYL22, CRT19]. Fast [BK18, LSW24, Mac23]. Feasible [Asa23]. field [FFO20]. filter [MCRV20, MPG18]. filter-based [MCRV20]. Filterbased [KLS18]. Filtering [MW20, Kar18, LY19a]. financial [AKW20, BCCI24, CWL23, FMU20, GP17, GOP18, KPY18, KJ20, Mor24]. Finding [CCCW23]. finish [ANL^{+}21]. Finite [BLvV21, BBP17]. finite-mixture [BBP17]. Fisher [For17]. Fisher-scoring [For17]. Fit [KALP24]. fitting [For17]. Fixed [WL21, BH24, May22]. Fixed-bandwidth [WL21]. Flexible [ACL19, Hau21, KJ20, AL17]. fMRI [DJF20, SAK21]. focus [PCP20]. Forecasting [BB24, HT23, HV21, KPY18, PL24, TWO24, BBS17, GOP18, GW17, PCP20, Sha17]. forecasts [Bay18]. form [PS19b]. Fourier [CS22]. Fractional [LS20, VS21]. framework [Kaw22]. Fraud [RPRH19]. free [Kiv23, Mac21a]. Frequency [BD22, PP23, ACG21, AJ21, Art24, BS18, CHM24, CS22, DKA17, LS20, LY19a, Mil18, Pol18, WCW18]. frontier [DFS20]. function [FR17, Ley23, Ric20]. function-on-scalar [FR17]. Functional [FKKP22, GM22, GSUC22, Sha17, ZPY^{+}24, AADN18, CKL^{+}23, FZF19, GFED22, KKW17, KOPS17, LW19, Mat20, MS17, SAK21]. Functions [LSY24, PZK22, LSBW20, ZDY22]. Future [LDA23]. Futures [KLMO22, Czu19]. Fuzzy [CRG23, Fer24, FGV21].

GARCH [ACG21, Asa23, BCFP23, CWL23, CK22, LCD21, ST24b, WCW18, WCD22]. GARCH-type [CWL23, LCD21]. Gaussian [KR22, Max20, MMS^{+}24, SC22]. GDP [RC22, ZYL22]. Gegenbauer [AMP20, PCP20]. gene [CZJ^{+}17]. gene-environment [CZJ^{+}17]. general [CR20, HT23]. Generalised [CKP21, OFG22]. Generalized [BCJC17, LSW24, SM20, AMK22, ADGG20, AMP20, BBP17, BK18, Egu18, KT19, PCP20, dVC18]. generative [HPZ22]. Global [ABC24, GOP18]. GMM [AR24, Hay19, HS20]. Goodness [KALP24]. Goodness-of-Fit [KALP24]. Google [BD22]. governance [GM21]. Gradient [AMK22]. Granger [AS19, SM19]. graphic [SV18]. graphical [GW17]. graphs [BB21]. Group [BS21]. grouped [HSS18]. grouping [ZDY22]. Growth [OFG22, HSS18]. guaranteed [ZMG18].

Hamiltonian [KC21]. Harris [AM19]. Hazard [ZY24]. health [BFS18]. heavy [BKP18, ES18]. heavy-tailed [BKP18, ES18]. HESSIAN [MMP21]. Heterogeneity [Mul18, DCVB22, Mar17, STI20, VS17]. heterogeneous [CRT19]. Heteroscedastic [CY21, PBMS20, GHK18, LDL21]. heteroscedasticity [LY19b, ySL21]. heteroscedasticity-consistent [ySL21]. heteroscedasticity-non-consistent [ySL21]. Heteroskedastic [SM20]. heteroskedasticity [LN17]. hidden [KLS18, LS21]. hierarchical [Col20, DCVB22, HGFO20]. High [BD22, CK22, FR17, LBWM22, LDA23, PP23, ZZZ20, ZY24, BBS17, CHM24, CS22, HO18, LS20, LY19b, Mil18, Mor19, WCW18, dVC18]. High-Dimensional [LBWM22, LDA23, CK22, FR17, ZZZ20, BBS17, HO18, LY19b]. High-Frequency [BD22, CHM24, CS22, LS20, Mil18, WCW18]. Higher [Mut18, LEP18]. Higher-order [Mut18]. highly [PS19a]. Hilbert [GRC17]. history [Fer24]. Homoscedastic [CY21]. horseshoe [FY19]. hyperplanes [Flo18]. hyperspectral [FZF19]. Hypothesis [CW23, Kir20, GJ19].

identifiability [DKA17]. Identification [AR24, Max20, Jia17, KPY18]. Identifying [CZJ^{+}17, Hay19]. identity [Ley23]. Imaging [ZPY^{+}24]. Imbalanced [LYBH21]. Imbalanced-Spatial-Structured [LYBH21]. Implicit [Smi23]. Improved [ES18, KR22, BS19]. Improving [DRW19]. impulse [LSBW20]. imputed [Cha23]. income [TB20, ZZ19]. Incorporating [VS17]. independent [Max20]. Index [TWO24, KPY18, Mor24]. indicator [FMU20]. indirect [Cre17, GM18, Mél22]. individual [For17]. Industrial [GOS24]. inequalities [KT19]. Inference [KR22, Lux22, MW23, ADG24, Bin18, CdL20, Cre17, DRW19, DF19, GM18, Kiv23, KC21, LEB17, NABM23, PB22, QD19, Ron20, SG17, WCD22]. inferences [GSSL20]. Infinitely [ZHO24]. Inflated [LSY24]. Inflation [BD22]. Influential [HPW23]. Information [RAC^{+}24, KH19, Lam18]. Inner [CLNP23]. Instrument [Kiv23]. Instrument-free [Kiv23]. Instrumental [CF21]. instruments [LEP18]. integer [CCW24]. Integrated [dZBMRV24]. integration [LEB17]. interactions [CZJ^{+}17]. interactive [BH24, NW19]. interest [Czu19]. intersection [Flo18]. interval [SVG19]. Intraday [DGK22, Sha17]. Introduction [GKD19]. Invariant [ZZZ20, ZZ19]. irregularly [ASJLZ17]. Issue [JS20, AKW20, BHL^{+}17, COSW17, CK21b, CK21c, DHW20, FKKP22, Gui18, GR23, HKLW18, HILM17, IL22, KOPS17, MPSS18]. issues [CK23, Kiv20]. Iterated [BB21]. Iterative [De 20].

January [Ano17d, Ano18b, Ano19g, Ano20h, Ano21h, Ano22e, Ano23e, Ano24h]. Joint [SM19, LSBW20]. July [Ano17c, Ano18a, Ano19e, Ano20e, Ano21f, Ano22g, Ano23h, Ano24e]. Jump [CK21a]. Jump-preserving [CK21a].

Kalman [MPG18]. Kernel [CKP21, CR20, FH19]. Kernel-based [CKP21]. kernels [FH21]. Knitting [BD22].

lag [Ang20, CR20]. lag-window [CR20]. lagged [De 20]. Lagrange [CA17]. Laplace [dZBMRV24, CMR18]. Large [BB24, Hau21, Li24, AJ21, HK22, HT23]. large-scale [HK22]. Lasso [BS21, Mac21a, Mac21b, MS17]. latent [FM17, For17, SSL23]. Learning [AM23, XY22, LY19a]. Least [CKP21, DRW19]. left [NABM23]. left-truncation [NABM23]. level [GG23, KPTV21, RC22]. Levels [RAC^{+}24]. leverage [CS22, KO20, PCP20]. Likelihood [CY21, WCD22, BP22, BCJC17, KO17, KH19, MCRV20, MOCB21, ST24b]. limit [PPDZ18, dVC18]. limited [Pol18]. limiting [Jia17]. line [GF19].

Linear [BB24, GM22, LSW24, MW20, AADN18, BBP17, Bla22, CJG24, CCW24, Egu18, FFO20, FH21, GKO18, GW17, KT19, Mac23, May22, PS19a, ySL21]. Lines [LDA23]. links [HVBJ17]. Local [GEH19, FFO20, FH21, May22, AO17]. locally [Kaw22]. located [ASJLZ17]. location [AMK22, SV18, TB17]. location-scale [SV18]. log [ADGG20, dVC18]. log-generalized [ADGG20, dVC18]. Logistic [PL24]. Long [BCR24, dBCR21, Art24, AMP20, GEH19, LS19, PCP20, WL21]. Long-Memory [dBCR21]. Longitudinal [LSY24, DJF20]. look [TB20]. low [Art24, FKM24, Mil18]. low-frequency [Mil18]. low-rank [FKM24].

M [BCJC17, SSL23]. Machine [LYBH21, XY22]. Macro [BB24, Mor24]. macro-financial [Mor24]. Macro-variables [BB24]. Magnetic [ZPY^{+}24]. maker [PPDZ18]. management [CWL23, GM21, MPSS18]. many [LEP18]. Marginal [PL24, Col20, GGHQ23, LEB17]. Market [ABC24, CG19, FMU20, PPDZ18, RSV20]. market-specific [CG19]. markets [Czu19, LS20, Mac21a, PPDZ18]. Markov [AMK22, ELS23, FS22, Kar18, KLS18, LS21, Lux22, NABM23, PS24, WCD22, YW23]. Markov-Switching [Lux22, PS24, AMK22, Kar18]. Markowitz [LBWM22]. martingale [GJ19, LDL21]. Matrices [Li24, BBS17, Mor19]. Matrix [GT23, Hir23, BV24, BCCI24]. Matrix-Variate [GT23]. matter [Sha17]. Max [ZHO24, GKO18]. Max-Infinitely [ZHO24]. max-linear [GKO18]. maximum [Flo18, KH19, MCRV20, ST24b]. MCMC [VQKS24]. Mean [KSS23, LBWM22, CK24, GFED22, HKTZ19]. Mean-Variance [LBWM22]. Means [Fer24]. measured [MZ22]. measurement [ZMG18]. Measures [DGK22, CJG24, ES18, KO20, STW18]. measuring [SFdB17]. mediation [GP23]. medical [GF19]. Memory [BCR24, dBCR21, Art24, AMP20, GEH19, LS19, PCP20, WL21]. Meta [ÖA17]. Meta-analytic [ÖA17]. Method [WW22, BKH22, CCW24, FH21, HK22, MMP21, NW19, Yos18]. Methodological [CK23]. Methods [BB24, CCCW23, Lux22, Beu23, COSW17, GRC17, GF19, HKLW18, Kiv20, LDL21, Mac23, ST24a]. Microeconometric [Kiv20]. MIDAS [ACG21, GQ19]. mild [Kiv23]. missing [Bin18, CdL20, HP21, Tan24]. Misspecification [BBP17, CdL20]. misspecified [Col20]. Mixed [SVG19, AJ21, CJG24, DKA17, HV21, LY19a, SZH19]. mixed-effects [CJG24]. mixed-frequency [AJ21, LY19a]. Mixing [XY22]. Mixture [CY21, Hau21, IL22, KR22, BBP17, DCVB22, FS22, HVBJ17, HILM17, MZ22, MBM17, SC22, ZDY22]. Mixtures [OFG22, XY22, CNBN22, GP17]. MLEs [BCR24]. mode [AMS23, FGV21]. Model [COySS23, Egu18, GT23, GM21, HPW23, KR22, Kiv20, LS19, MW20, SMSC21, ZYL22, ZY24, AL17, ADGG20, ACG21, AM19, BBS17, BCCI24, CdL20, CJG24, CCW24, DCVB22, ELS23, HT23, HKTZ19, HGFO20, KO17, KH19, KC21, MZ22, Mél22, SOLP17, ST24b, SSL23, SZH19, Tan24, Yos18, ZMG18, ZDY22].

Model-Based [MW20]. Modeling [ABC24, CG19, FFO20, PZK22, BFS18, GKD19, HPZ22, Tan24]. Modelling [CIO19, Fok24, OFG22]. Models [Asa23, BCR24, BLvV21, CDH23, CF21, Hau21, IL22, KLMO22, LSW24, LDA23, Lux22, OFG22, PS24, SM20, ACL19, AMK22, AADN18, Ang20, AGE17, AMP20, BBP17, dZBMRV24, Beu23, BM22, BCJC17, BKH22, CLNP23, CJG24, CRT19, CA17, Cha22, CWL23, CK21a, Col20, De 20, DF21, Egu18, FM17, For17, GKO18, GM21, GW17, HK22, HL23, HVBJ17, Hay19, HV21, HILM17, HS20, HP21, Hu21, HLU20, Kar18, KT19, KLS18, KJ20, Lam18, LY19a, LS19, LDL21, LCD21, LS21, LN17, LSBW20, Mac23, Mac21b, MCRV20, MPG18, Mat20, May22, MTV17, MOCB21, Mut18, NABM23, NW19, PS19a, PCP20, PB22, PBMS20, PS19b, Ric20, SC22, SG17, SM19, SV18, TB17, VS17, WFSJ23, WCW18, WCD22, Yos18, ZZ19]. moderation [GP23]. modification [Döh18]. moments [BKH22]. Monitoring [RPRH19]. monotonic [Dem21]. Monte [KC21, BLvV21, LCD21, WCD22]. monthly [HKTZ19]. moving [NW19, PS19b]. Multi [BD22, BSM23, ZPY^{+}24, HGFO20, NABM23, PRV17, RC22]. Multi-Annual [BD22]. multi-level [RC22]. Multi-objective [BSM23]. multi-sample [PRV17]. multi-state [NABM23]. Multi-Subject [ZPY^{+}24]. multi-trial [HGFO20]. Multifractal [Lux22]. Multinomial [MS17, Col20, HVBJ17]. Multiple [KO20, OFG22, Max20, MTV17, SM19]. Multiple-block [KO20]. multiplicative [ST24b]. multiplier [CA17]. multiply [Cha23]. multiply-imputed [Cha23]. Multivariate [COySS23, CY21, Fok24, HPZ22, MMP21, MW20, VQKS24, AFKL21, ACL19, BV24, BCFP23, GW17, JLM24, MCNO17, MGF18, MMS^{+}24, SS23, STI20].

Near [BB24]. Near-equivalence [BB24]. Nearly [AR24]. Nearly-Weak [AR24]. nested [dZBMRV24, Bla22]. nets [Cre17]. network [SM19]. Networks [BCFP23, BCCI24, HPZ22]. Neural [Cre17, HPZ22]. Neuroimaging [JS20]. noise [CR20, GKO18]. noisy [DFS20]. Non [DKA17, Beu23, DF21, FFO20, GOP18, LEP18, NABM23, SC22, ySL21]. non- [Beu23]. non-Gaussian [SC22]. Non-identifiability [DKA17]. non-linear [FFO20]. non-Markov [NABM23]. non-normal [LEP18]. non-stationary [DF21]. non-synchronous [GOP18]. noncausal [HV21]. nonconstant [Mul18]. nonignorable [Tan24]. Nonlinear [Lux22, OFG22, PL24, ASJLZ17, Cha22, CK21a]. Nonparametric [CF21, FZF19, FH19, LEB17, CLNP23, CNBN22, GD22, SV18, ZDY22]. Normal [CY21, AFKL21, GP17, Ley23, LEP18]. normality [PV17, ST24b]. Note [BS21]. novel [SFdB17]. Nowcasting [ZYL22, AJ21]. null [MGF18]. Number [ZYL22, LS21]. Numerical [CCCW23, HK22].

obesity [CKL^{+}23]. objective [BSM23]. Objects [PZK22]. observables [HP21]. observational [LEB17]. Observations [HPW23, BI19, Egu18]. occupation [NABM23]. October [Ano17a, Ano18c, Ano19h, Ano20g, Ano21g, Ano22f, Ano23g, Ano24g]. OLS [GQ19]. On-line [GF19]. one [ADG24]. one-sided [ADG24]. online [GW17]. open [Czu19]. operator [CR20]. Optimal [CNBN22, LSW24, AO17, CCCW23]. Optimality [BLvV21]. optimisation [BSM23]. Optimization [LBWM22, ST24a]. option [Mac21a, Mac21b, ZMG18]. Oracle [KT19]. order [LS19, LEP18, Mut18, PPDZ18]. ordered [BBP17]. ordering [BS19]. ordinal [MZ22, STI20, TB17]. orthogonal [GG23]. outcomes [STI20, WFSJ23]. Outlier [RAC^{+}24]. outliers [CdL20]. over-identifying [Hay19]. overparameterization [AFKL21]. oversampling [Pol18]. Overview [Smi23].

Pages [Ano17b, Ano17d, Ano17c, Ano17a, Ano18b, Ano18a, Ano18c, Ano19f, Ano19g, Ano19e, Ano19h, Ano20f, Ano20h, Ano20e, Ano20g, Ano21e, Ano21h, Ano21f, Ano21g, Ano22e, Ano22g, Ano22f, Ano23f, Ano23e, Ano23h, Ano23g, Ano24f, Ano24h, Ano24e, Ano24g]. Panel [Asa23, AÖ21, BKH22, CRT19, Hay19, HLU20, Kiv20, LW19, Mac21b, NW19, PB22, RC22, WFSJ23, ZZ19]. Panels [BB24, De 20, ÖA17]. Parameter [Hau21, LW19, GQ19, HK22, Kir20]. parameters [CCW24, Kir20]. parametric [BP22, HVBJ17]. Paretian [Pao17]. parsimonious [AL17]. parsimony [FY19]. partial [Jia17, LW19]. Partially [ADG24, GG23]. Particle [LY19a, Kar18, MCRV20]. particulate [Sha17]. payoffs [RSV20]. peak [GF19]. penalization [MS17]. penalized [Bay18, KV24]. penalties [KV24]. Penalty [ZY24]. performance [KP17]. periodic [LS19, ZP21]. Periodicity [DGK22]. persistent [ADG24]. Perspective [AM23, BCFP23]. Phillips [KPTV21]. plot [BSM23]. Point [SWS22, BHKX17, ELS23, GHK18]. Points [ABC24]. polygenic [CKL^{+}23]. polynomial [GQ19]. portfolio [GP17, GSSL20, HW24, ST24a]. portfolios [HT23, PS23]. positive [FH19]. possibly [Col20]. posteriors [Jia17]. potentials [FFO20]. power [May22]. predetermined [PB22]. Predict [BD22]. Prediction [KKW17, MW20, WW22]. predictions [STW18]. Predictive [HPW23]. predictor [FZF19]. predictors [Mil18]. Preliminary [PRV17]. premia [PS23, RSV20]. presence [Max20, MTV17]. preserving [AMS23, CK21a]. Price [TWO24, SVG19]. pricing [Mac21b, PS23]. principal [PRV17]. Prior [HPW23, FY19]. Priors [Hau21, Ber20, FHT20, GP23]. Private [KALP24]. probabilistic [ANL^{+}21]. Probabilities [PS24, NABM23]. Probability [LSW24, PZK22]. Problem [MW20, MW23, Dem21]. procedure [Döh18]. Process [SWS22, AM19, CK22, FM17, YW23]. Processes [dBCR21, VQKS24, ZHO24, ADG24, DF19, KKW17, Pol18, PV17]. profiling [GQ19]. prognosis [CZJ^{+}17]. proof [Mél22]. Properties [MCRV20, QD19, Mél22]. Proportions [XY22]. pseudo [GM21]. pseudo-models [GM21].

Quadratic [Mat20]. Quantile [AGMR18, LSY24, Mac21a, Mac21b, ADGG20, BSTS22, Bay18, HKLW18, Mul18, OSSRK19, WCW18, dVC18]. quasi [FHT20, Jia17]. quasi-Bayesian [FHT20]. quasi-posteriors [Jia17].

Rage [KSS23]. random [BBP17, Bin18, CRG23, HP21, JLM24, NABM23]. Randomized [COySS23, SM20]. rank [ANL^{+}21, AÖ21, Bin18, FKM24, FMU20, ÖA17]. rat [FFO20]. Ratio [CY21, BS19]. Realized [AMP20, DGK22, TWO24, BBS17, CWL23, CIO19, GSSL20, KO20, SOLP17, SVG19]. Reconstruction [KC23]. Recursive [WW22, GKO18]. reduced [ANL^{+}21]. reducing [GEH19]. Reduction [BB24, KPY18, MCNO17, ST24a]. Regimes [OFG22, LW19]. Regression [Ber20, CDH23, GM22, KC23, KSS23, ASJLZ17, BB21, BSTS22, BCCI24, Bla22, FR17, FZF19, FH21, GF19, HKLW18, Hu21, LW19, Mac23, Mat20, MS17, Mul18, MMS^{+}24, OSSRK19, PS19a, ySL21, SV18, TB17, WCW18, Yos18]. Regressions [PL24, XY22, Bay18, GQ19, HK22, LY19b]. regressive [SZH19]. regressor [Kiv23]. Regressors [CF21, PS19a]. regularity [Kiv23]. regularization [DFS20, HW24]. Regularized [Mor19]. regularly [GKO18]. Relevance [MW23]. reliable [BK18]. Reproducibility [KR22]. Research [LDA23]. Resonance [ZPY^{+}24]. Response [CF21, COySS23, GG23, LSBW20, Mat20, STI20, YW23]. resting [DJF20]. resting-state [DJF20]. restricted [DF21]. restriction [Hay19]. results [PS19b]. Retrospective [KC23]. Returns [TWO24]. Review [AM23, CJG24, Hir23, KSS23, Li24, Beu23, LN17, LSBW20]. right [NABM23]. right-censoring [NABM23]. Risk [Bay18, CK22, GD22, ST24a, BCFP23, CWL23, CKL^{+}23, DCVB22, ES18, GM21, HT23, MPSS18, PS23, RSV20, ZMG18]. risks [GM18]. Robust [BH24, CDH23, DFS20, GP17, GJ19, Hir23, KV24, RPRH19, CZJ^{+}17, MPG18, Mil18, Ron20, SVG19]. Robustness [CdL20, HPW23, Mar17]. Role [BCR24]. root [CHM24, MTV17, MGF18, ZP21]. Rosenblatt [SS23]. Rosenblatt-type [SS23]. rules [Cha23].

S&P [HO18]. Sample [BLvV21, KC23, ZZZ20, PRV17]. sampling [AADN18, CHM24]. scalable [GW17]. scalar [FR17]. Scale [ZZZ20, AMK22, HK22, SV18]. Scale-Invariant [ZZZ20]. scales [MZ22]. scanned [KP17]. scatter [BV24]. Score [BLvV21, GM22, CKL^{+}23, Flo18, Ley23]. Score-Driven [BLvV21]. scoring [For17]. SDB [MBM17]. Seasonal [dBCR21, HKTZ19, ZP21]. Seasonality [PP23, HKTZ19]. sectional [AÖ21, JM21, RC22]. sector [HW24]. sector-specific [HW24]. seemingly [HK22]. segmentation [CK22, CK24]. selecting [ZDY22]. Selection [Col20, Ang20, AO17, BP22, CJG24, GW17, HW24, Kiv20, KH19, LS19]. Selective [Li24, LDA23]. semantic [MZ22]. Semi [AM23, BP22]. semi-parametric [BP22]. Semi-Supervised [AM23]. Semiparametric [BFS18, CdL20, PL24, WS18, XY22, Yos18, ADG24, ASJLZ17, Beu23, BCCI24, BCJC17, HKLW18, HLU20, Mor19, OSSRK19, SZH19, Tan24]. Sensitive [COySS23]. separable [GRC17]. Separating [TB17]. Sequential [LCD21, WCD22]. Series [BCR24, BS21, Fok24, GT23, Hir23, PL24, PP23, RPRH19, VQKS24, ACL19, ASJLZ17, Art24, BM22, BHL^{+}17, BCJC17, Cha22, CK21a, ELS23, FM17, GHK18, GF19, HKTZ19, HPZ22, LEB17, MCNO17, Mil18, Sha17, SAK21]. service [BFS18]. sets [CRG23, SMSC21]. setting [FGV21]. shape [AMK22, WS18]. shifting [HKTZ19]. shocks [Max20]. Short [BCR24, De 20]. shortfall [BKP18, GGHQ23, HT23]. sided [ADG24]. Sign [BI19, BV24, KT19]. signal [AGE17]. signals [HGFO20]. signed [Bin18]. signed-rank [Bin18]. Significance [RAC^{+}24]. Simple [Mil18, ZZZ20]. simulated [Flo18, KO17, MCRV20]. Simulation [AJ21, ZHO24, KP17, Lam18]. Simultaneous [GFED22, GW17]. single [KC21, PBMS20]. Singular [AGE17]. skew [AFKL21, MBM17]. skew- [MBM17]. skew-normal [AFKL21]. skewing [FH21]. slab [CLNP23]. smooth [ST24b]. smoothers [ANL^{+}21]. smoothing [AJ21, LY19a, THS18]. Solutions [MW20]. Some [BLvV21, BV24, Cha22, May22]. Space [KLMO22, Lux22, Kir20, LY19a, MPG18]. spaces [GRC17]. sparsity [Hu21]. Spatial [LYBH21, ZPY^{+}24, ASJLZ17, SZH19]. Spatial-Temporal [ZPY^{+}24]. Spatially [Hu21]. spatio [ADN18, MOCB21]. spatio-temporal [ADN18, MOCB21]. Special [BHL^{+}17, COSW17, CK23, FKKP22, Gui18, GR23, HKLW18, HILM17, IL22, JS20, KOPS17, MPSS18, DHW20, GKD19]. specific [CG19, HW24]. specification [Kiv20, Mil18, Tan24]. specifications [Bla22]. Spectral [OP24, VQKS24, CR20, DJF20, FFO20, Kaw22]. Spectrally [LBWM22]. Spectrally-Corrected [LBWM22]. Spectrum [AGE17]. spike [CLNP23]. spillovers [BCFP23]. spline [ANL^{+}21, KV24, Ric20]. splines [THS18]. split [BSM23]. split-plot [BSM23]. Spot [CMR18]. spread [JM21]. spreading [BCCI24]. Spurious [JM21]. square [CHM24]. square-root [CHM24]. squared [QD19]. Squares [CKP21, DRW19]. Stable [SM20, Pao17]. stage [CRT19, KH19, Mul18]. Standard [GEH19]. State [KPTV21, KLMO22, Lux22, PS24, DJF20, LY19a, MPG18, NABM23]. State-Dependent [PS24]. State-level [KPTV21]. State-Space [KLMO22, LY19a]. states [LS21]. stationarity [MGF18]. Stationary [VQKS24, DF21, Kaw22, PV17]. Statistic [CW23]. Statistical [AM23, GSSL20, MW23, NABM23]. Statistics [CK21b, CK21c, CK23, Cha23, COSW17, Gui18, Mut18, CRG23, KVC17]. Stochastic [CS22, KLMO22, Pol18, ZYL22, AM19, AMP20, dZBMRV24, GOP18, Kar18, KC21, KLS18, MCRV20, MMP21, MGF18, PS23, PCP20, SG17, SOLP17, AGE17]. Stock [TWO24, SVG19]. storage [KO17]. storms [BHKX17]. STOXX [CG19]. strategy [AO17]. stratification [CNBN22]. stratified [PBMS20]. streams [ADN18]. strict [May22]. Structural [BS21, LN17, AÖ21, KPY18, LS21, Max20, MTV17]. structure [Hay24, HS20, SAK21, Yos18]. Structured [LYBH21]. structures [ZDY22].

Student [PCP20]. Student- [PCP20]. STUR [MGF18]. Subject [ZPY^{+}24, NABM23]. Subjective [STI20]. Subsampling [LSW24, VQKS24]. subsets [KP17]. subspace [AS19]. suitable [Ric20]. sums [JLM24]. Supervised [AM23, MCNO17]. Support [LYBH21]. supported [FH19]. surface [ANL^{+}21, GG23]. Survey [LDA23]. survival [CNBN22, VS17]. swap [RSV20]. Switching [Lux22, PS24, AMK22, Kar18, WCD22]. synchronous [GOP18]. synthetic [BK18]. systems [DKA17, PBMS20].

Tail [GKO18, ADGG20, CWL23, Kir20, dVC18]. tailed [BKP18, ES18]. tale [Ley23]. targeted [KPY18]. Targeting [Asa23]. techniques [GEH19]. temperature [HKTZ19]. Temporal [ZPY^{+}24, ADN18, Cha22, CIO19, MOCB21]. term [HS20]. Test [CY21, GM22, ZZZ20, AÖ21, BBP17, BS19, CA17, CR20, Hay19, MGF18, PRV17, PV17, Tan24, ZP21]. Testing [AS19, Bla22, CW23, LY19b, MGF18, BCJC17, FH19, Kir20, MTV17, Pao17]. Tests [KALP24, BV24, BI19, Col20, KP17, May22, Mil18, ÖA17, WL21]. theoretic [Lam18]. theoretical [DHW20]. theory [PS23]. three [GG23]. three-level [GG23]. threshold [dZBMRV24]. Tikhonov [DFS20]. Time [BCR24, BS21, BM22, Fok24, GT23, Hau21, Hir23, PL24, PP23, PS24, RPRH19, VQKS24, ACL19, ASJLZ17, Art24, BHL^{+}17, BCJC17, Cha22, CK21a, Czu19, ELS23, FM17, GHK18, GF19, HK22, HSS18, HPZ22, KLS18, LS20, LEB17, MCNO17, STW18, Sha17, ST24b, SAK21, ZZ19]. time-invariant [ZZ19]. time-series [ASJLZ17, Cha22, HPZ22]. time-to-event [STW18]. Time-Varying [PS24, Hau21, Czu19, HK22, HSS18, LS20, ST24b]. Tobit [HL23]. tool [TB20]. topic [GKD19]. tractable [AL17]. trading [Czu19]. traffic [KKW17]. transform [CMR18]. transformation [SS23]. Transformed [GT23]. transforms [BM22]. Transition [PS24, NABM23, ST24b]. Treasury [THS18]. treatment [AGMR18, WFSJ23]. trend [BHKX17]. trending [ADG24]. Trends [BD22, GOP18]. trial [HGFO20]. tropical [BHKX17]. truncation [FH19, NABM23]. Tuning [HPW23]. Turning [ABC24]. Two [GT23, ZZZ20, CRT19, FGV21, KH19, Mul18, PBMS20, ZMG18]. two-decrement [ZMG18]. two-mode [FGV21]. Two-Sample [ZZZ20]. two-stage [CRT19, KH19, Mul18]. Two-Way [GT23, PBMS20]. Type [ZY24, CWL23, DF19, GGHQ23, LCD21, SS23].

UK [KPY18]. unit [MTV17, MGF18, ZP21]. Univariate [CK24, PS19b]. Unknown [ZYL22, LS21]. unobserved [DCVB22, VS17]. unrelated [HK22]. updating [Sha17]. use [LEP18]. Using [KLMO22, ySL21, TWO24, ZYL22, AFKL21, ANL^{+}21, Bay18, BM22, CZJ^{+}17, CKL^{+}23, CMR18, FFO20, FH21, GP23, GOP18, HVBJ17, KPY18, KR22, KC21, LDL21, MMP21, ST24a, WCW18, WCD22]. v [BM22]. v-transforms [BM22]. validation [Lam18, SMSC21]. Valuation [GM21, ZMG18]. Value [Bay18, CK22, GD22, HT23]. Value-at-Risk [Bay18, CK22, GD22, HT23]. Vandermonde [FKM24]. VAR [LSBW20]. Variable [CF21, De 20, GW17, SSL23]. Variables [KALP24, BB24, GKO18, KP17, MZ22, TB20]. Variance [Asa23, LBWM22, RSV20, SVG19]. Variance-Targeting [Asa23]. variances [CIO19, ySL21]. Variate [GT23]. variation [CJG24, ZP21]. VARMA [DKA17, Mél22]. VARs [AJ21]. Varying [PS24, XY22, CK21a, Czu19, GKO18, HK22, HSS18, Hau21, Hu21, LS20, ST24b, SZH19]. varying-coefficient [CK21a, SZH19]. Vector [LYBH21, VQKS24, CA17, FY19, LN17, PS19b]. versus [GEH19]. very [ES18]. via [CNBN22, FH21, GQ19, LCD21, SMSC21]. Vine [SC22, ACL19]. vines [BM22, CIO19]. Visualizing [HO18]. VMA [DKA17]. Volatility [BLvV21, CKP21, DGK22, GOP18, KLMO22, TWO24, WW22, ZYL22, ACG21, AM19, AMP20, dZBMRV24, CG19, CMR18, Czu19, GW17, Kar18, KC21, KLS18, LS20, LN17, MCRV20, MMP21, PCP20, SG17, SOLP17, SVG19, AGE17]. volume [Czu19].

wage [KPTV21]. Wang [ES18]. wavelet [LDL21]. wavelets [MS17]. Way [GT23, PBMS20]. Weak [AR24]. Weibull [ADGG20, VS17, dVC18]. Weibull-tail [ADGG20]. weighted [DRW19]. weights [GSSL20]. Weissman [GGHQ23]. Weissman-type [GGHQ23]. white [CR20]. Whittle [AO17, GEH19]. wide [PV17]. window [CR20]. Wishart [KLMO22]. within [ACG21, BS18].

X [ACG21].

Yekutieli [Döh18]. yield [THS18].

Zellner [HPW23]. Zero [LSY24]. Zero-Inflated [LSY24].

References

Ahmed:2018:BFL

[AADN18] M. S. Ahmed, M. K. Attouch, and S. Dabo-Niang. Binary functional linear models under choice-based sampling. Econometrics and Statistics, 7(??):134–152, July 2018. CODEN ???? ISSN

            2452-3062. URL http://www.sciencedirect.com/science/
            article/pii/S245230621730062X.

Ahelegbey:2024:MTP

[ABC24] Daniel Felix Ahelegbey, Monica Billio, and Roberto Casarin. Modeling turning points in the global equity market. Econometrics and Statistics, 30(??):60–75, April 2024. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001192.

Amendola:2021:CFV

[ACG21] Alessandra Amendola, Vincenzo Candila, and Giampiero M. Gallo. Choosing the frequency of volatility components within the double asymmetric GARCH–MIDAS–X model. Econometrics and Statistics, 20(??):12–28, October 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221000071.

Acar:2019:FDV

[ACL19] Elif F. Acar, Claudia Czado, and Martin Lysy. Flexible dynamic vine copula models for multivariate time series data. Econometrics and Statistics, 12(??):181–197, October 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306219300206.

Abadir:2024:POS

[ADG24] Karim M. Abadir, Walter Distaso, and Liudas Giraitis. Partially one-sided semiparametric inference for trending persistent and antipersistent processes. Econometrics and Statistics, 30(??):1–14, April 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001611.

Albert:2020:EQE

[ADGG20] Clément Albert, Anne Dutfoy, Laurent Gardes, and Stéphane Girard. An extreme quantile estimator for the log-generalized Weibull-tail model. Econometrics and Statistics, 13(??):137– 174, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300097.

Amiri:2018:DES

[ADN18] Aboubacar Amiri and Sophie Dabo-Niang. Density estimation over spatio-temporal data streams. Econometrics and Statistics, 5(??):148–170, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300734.

Abe:2021:EAU

[AFKL21] Toshihiro Abe, Hironori Fujisawa, Takayuki Kawashima, and Christophe Ley. EM algorithm using overparameterization for the multivariate skew-normal distribution. Econometrics and Statistics, 19(??):151–168, July 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000332.

Arteche:2017:SSA

[AGE17] Josu Arteche and Javier García-Enríquez. Singular Spectrum Analysis for signal extraction in Stochastic Volatility models. Econometrics and Statistics, 1(??):85–98, January 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300156.

Alejo:2018:QCT

[AGMR18] Javier Alejo, Antonio F. Galvao, and Gabriel Montes-Rojas. Quantile continuous treatment effects. Econometrics and Statistics, 8(??):13–36, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300928.

Ankargren:2021:SSN

[AJ21] Sebastian Ankargren and Paulina Jonéus. Simulation smoothing for nowcasting with large mixed-frequency VARs. Econometrics and Statistics, 19(??):97–113, July 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306220300538.

Ando:2020:CAC

[AKW20] Tomohiro Ando, Erricos Kontoghiorghes, and Peter Winker. CFEnetwork: the annals of computational and financial econometrics, 5^{th} issue. Econometrics and Statistics, 13 (??):1, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300656.

Abe:2017:TPF

[AL17] Toshihiro Abe and Christophe Ley. A tractable, parsimonious and flexible model for cylindrical data, with applications. Econometrics and Statistics, 4(??):91–104, October 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300168.

Anzarut:2019:HPM

[AM19] Michelle Anzarut and Ramsés H. Mena. A Harris process to model stochastic volatility. Econometrics and Statistics, 10(??):151–169, April 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300030.

Ahfock:2023:SSL

[AM23] Daniel Ahfock and Geoffrey J. McLachlan. Semi-supervised learning of classifiers from a statistical perspective: a brief review. Econometrics and Statistics, 26(??):124–138, April 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306222000296.

Adam:2022:GBM

[AMK22] Timo Adam, Andreas Mayr, and Thomas Kneib. Gradient boosting in Markov-switching generalized additive models for location, scale, and shape. Econometrics and Statistics, 22(??):3–16, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000502.

Asai:2020:RSV

[AMP20] Manabu Asai, Michael McAleer, and Shelton Peiris. Realized stochastic volatility models with generalized Gegenbauer long memory. Econometrics and Statistics, 16(??): 42–54, October 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300048.

Abe:2023:BEM

[AMS23] Toshihiro Abe, Yoichi Miyata, and Takayuki Shiohama. Bayesian estimation for mode and anti-mode preserving circular distributions. Econometrics and Statistics, 27(??):136–160, July 2023. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306221000344.

Angelini:2020:BLS

[Ang20] Giovanni Angelini. Bootstrap lag selection in DSGE models with expectations correction. Econometrics and Statistics, 14(??):38–48, April 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300874.

Arnqvist:2021:ESF

[ANL^{+}21] Natalya Pya Arnqvist, Blaise Ngendangenzwa, Eric Lindahl, Leif Nilsson, and Jun Yu. Efficient surface finish defect detection using reduced rank spline smoothers and probabilistic classifiers. Econometrics and Statistics, 18(??):89–105, April 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300514.

Anonymous:2017:PO

[Ano17a] Anonymous. Pages 1-130 (October 2017). Econometrics and Statistics, 4(??):1–130, October 2017. CODEN ???? ISSN 2452-3062.

Anonymous:2017:PA

[Ano17b] Anonymous. Pages 1-148 (April 2017). Econometrics and Statistics, 2(??):1–148, April 2017. CODEN ???? ISSN 2452-3062.

Anonymous:2017:PJb

[Ano17c] Anonymous. Pages 1-168 (July 2017). Econometrics and Statistics, 3(??):1–168, July 2017. CODEN ???? ISSN 2452-3062.

Anonymous:2017:PJa

[Ano17d] Anonymous. Pages 1-200 (January 2017). Econometrics and Statistics, 1(??):1–200, January 2017. CODEN ???? ISSN 2452-3062.

Anonymous:2018:PJb

[Ano18a] Anonymous. Pages 1-164 (July 2018). Econometrics and Statistics, 7(??):1–164, July 2018. CODEN ???? ISSN 2452-3062.

Anonymous:2018:PJa

[Ano18b] Anonymous. Pages 1-188 (January 2018). Econometrics and Statistics, 5(??):1–188, January 2018. CODEN ???? ISSN 2452-3062.

Anonymous:2018:PO

[Ano18c] Anonymous. Pages 1-250 (October 2018). Econometrics and Statistics, 8(??):1–250, October 2018. CODEN ???? ISSN 2452-3062.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial board. Econometrics and Statistics, 9(??):ii, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621830100X.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Econometrics and Statistics, 10 (??):ii, April 2019. CODEN ???? ISSN 2452-3062. URL https://

www.sciencedirect.com/science/article/pii/S2452306219300164.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Econometrics and Statistics, 11 (??):ii, July 2019. CODEN ???? ISSN 2452-3062. URL https://

www.sciencedirect.com/science/article/pii/S2452306219300358.

Anonymous:2019:EB

[Ano19d] Anonymous. Editorial Board. Econometrics and Statistics, 12 (??):ii, October 2019. CODEN ???? ISSN 2452-3062. URL

https://www.sciencedirect.com/science/article/pii/S2452306219300553.

Anonymous:2019:PJb

[Ano19e] Anonymous. Pages 1-158 (July 2019). Econometrics and Statistics, 11(??):1–158, July 2019. CODEN ???? ISSN 2452-3062.

Anonymous:2019:PA

[Ano19f] Anonymous. Pages 1-170 (April 2019). Econometrics and Statistics, 10(??):1–170, April 2019. CODEN ???? ISSN 2452-3062.

Anonymous:2019:PJa

[Ano19g] Anonymous. Pages 1-170 (January 2019). Econometrics and Statistics, 9(??):1–170, January 2019. CODEN ???? ISSN 2452-3062.

Anonymous:2019:PO

[Ano19h] Anonymous. Pages 1-216 (October 2019). Econometrics and Statistics, 12(??):1–216, October 2019. CODEN ???? ISSN 2452-3062.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Econometrics and Statistics, 13(??):ii, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300101.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Econometrics and Statistics, 14(??):ii, April 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300277.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Econometrics and Statistics, 15(??):ii, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300629.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Econometrics and Statistics, 16(??):ii, October 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300770.

Anonymous:2020:PJb

[Ano20e] Anonymous. Pages 1-136 (July 2020). Econometrics and Statistics, 15(??):1–136, July 2020. CODEN ???? ISSN 2452-3062.

Anonymous:2020:PA

[Ano20f] Anonymous. Pages 1-158 (April 2020). Econometrics and Statistics, 14(??):1–158, April 2020. CODEN ???? ISSN 2452-3062.

Anonymous:2020:PO

[Ano20g] Anonymous. Pages 1-168 (October 2020). Econometrics and Statistics, 16(??):1–168, October 2020. CODEN ???? ISSN 2452-3062.

Anonymous:2020:PJa

[Ano20h] Anonymous. Pages 1-196 (January 2020). Econometrics and Statistics, 13(??):1–196, January 2020. CODEN ???? ISSN 2452-3062.

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Econometrics and Statistics, 17(??):ii, January 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622100023X.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Econometrics and Statistics, 18(??):ii, April 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000411.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Econometrics and Statistics, 19(??):ii, July 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000691.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Econometrics and Statistics, 20(??):ii, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000897.

Anonymous:2021:PA

[Ano21e] Anonymous. Pages 1–142 (April 2021). Econometrics and Statistics, 18(??):1–142, April 2021. CODEN ???? ISSN 2452-3062.

Anonymous:2021:PJb

[Ano21f] Anonymous. Pages 1–188 (July 2021). Econometrics and Statistics, 19(??):1–188, July 2021. CODEN ???? ISSN 2452-3062.

Anonymous:2021:PO

[Ano21g] Anonymous. Pages 1–202 (October 2021). Econometrics and Statistics, 20(??):1–202, October 2021. CODEN ???? ISSN 2452-3062.

Anonymous:2021:PJa

[Ano21h] Anonymous. Pages 1-172 (January 2021). Econometrics and Statistics, 17(??):1–172, January 2021. CODEN ???? ISSN 2452-3062.

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Econometrics and Statistics, 21(??):ii, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001544.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Econometrics and Statistics, 22(??):ii, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622200020X.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Econometrics and Statistics, 23(??):ii, July 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622200034X.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Econometrics and Statistics, 24(??):ii, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000922.

Anonymous:2022:PJa

[Ano22e] Anonymous. Pages 1–178 (January 2022). Econometrics and Statistics, 21(??):1–178, January 2022. CODEN ???? ISSN 2452-3062.

Anonymous:2022:PO

[Ano22f] Anonymous. Pages 1–194 (October 2022). Econometrics and Statistics, 24(??):1–194, October 2022. CODEN ???? ISSN 2452-3062.

Anonymous:2022:PJb

[Ano22g] Anonymous. Pages 1–204 (July 2022). Econometrics and Statistics, 23(??):1–204, July 2022. CODEN ???? ISSN 2452-3062.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial Board. Econometrics and Statistics, 25(??):ii, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222001022.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial Board. Econometrics and Statistics, 26(??):ii, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306223000175.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial Board. Econometrics and Statistics, 27(??):ii, July 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622300045X.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial Board. Econometrics and Statistics, 28(??):ii, October 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306223000667.

Anonymous:2023:PJa

[Ano23e] Anonymous. Pages 1–134 (January 2023). Econometrics and Statistics, 25(??):1–134, January 2023. CODEN ???? ISSN 2452-3062.

Anonymous:2023:PAa

[Ano23f] Anonymous. Pages 1–160 (April 2023). Econometrics and Statistics, 26(??):1–160, April 2023. CODEN ???? ISSN 2452-3062.

Anonymous:2023:PO

[Ano23g] Anonymous. Pages 1–172 (October 2023). Econometrics and Statistics, 28(??):1–172, October 2023. CODEN ???? ISSN 2452-3062.

Anonymous:2023:PJb

[Ano23h] Anonymous. Pages 1–196 (July 2023). Econometrics and Statistics, 27(??):1–196, July 2023. CODEN ???? ISSN 2452-3062.

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial Board. Econometrics and Statistics, 29(??):ii, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306223000862.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial Board. Econometrics and Statistics, 30(??):ii, April 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306224000248.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial Board. Econometrics and Statistics, 31(??):ii, July 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306224000443.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial Board. Econometrics and Statistics, 32(??):ii, October 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306224000637.

Anonymous:2024:PJb

[Ano24e] Anonymous. Pages 1–130 (July 2024). Econometrics and Statistics, 31(??):1–130, July 2024. CODEN ???? ISSN 2452-3062.

Anonymous:2024:PAa

[Ano24f] Anonymous. Pages 1–132 (April 2024). Econometrics and Statistics, 30(??):1–132, April 2024. CODEN ???? ISSN 2452-3062.

Anonymous:2024:PO

[Ano24g] Anonymous. Pages 1–160 (October 2024). Econometrics and Statistics, 32(??):1–160, October 2024. CODEN ???? ISSN 2452-3062.

Anonymous:2024:PJa

[Ano24h] Anonymous. Pages 1–282 (January 2024). Econometrics and Statistics, 29(??):1–282, January 2024. CODEN ???? ISSN 2452-3062.

Arteche:2017:SOB

[AO17] Josu Arteche and Jesus Orbe. A strategy for optimal bandwidth selection in Local Whittle estimation. Econometrics and Statistics, 4(??):3–17, October 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300041.

Arsova:2021:PCR

[AÖ21] Antonia Arsova and Deniz Dilan Karaman Örsal. A panel cointegrating rank test with structural breaks and crosssectional dependence. Econometrics and Statistics, 17(??): 107–129, January 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300484.

Antoine:2024:GNW

[AR24] Bertille Antoine and Eric Renault. GMM with nearlyweak identification. Econometrics and Statistics, 30(??): 36–59, April 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001258.

Arteche:2024:BLM

[Art24] Josu Arteche. Bootstrapping long memory time series: Application in low frequency estimators. Econometrics and Statistics, 29(??):1–15, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000769.

Al-Sadoon:2019:TSG

[AS19] Majid M. Al-Sadoon. Testing subspace Granger causality. Econometrics and Statistics, 9(??):42–61, January 2019. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300709.

Asai:2023:FPG

[Asa23] Manabu Asai. Feasible panel GARCH models: Variance-targeting estimation and empirical application. Econometrics and Statistics, 25(??):23–38, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622200003X.

Al-Sulami:2017:ESN

[ASJLZ17] Dawlah Al-Sulami, Zhenyu Jiang, Zudi Lu, and Jun Zhu. Estimation for semiparametric nonlinear regression of irregularly located spatial time-series data. Econometrics and Statistics, 2(??):22–35, April 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300047.

Bayer:2018:CVR

[Bay18] Sebastian Bayer. Combining value-at-risk forecasts using penalized quantile regressions. Econometrics and Statistics, 8 (??):56–77, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300680.

Baranyi:2021:ICE

[BB21] Máté Baranyi and Marianna Bolla. Iterated conditional expectation algorithm on DAGs and regression graphs. Econometrics and Statistics, 20(??):131–152, October 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300496.

Barbarino:2024:FNE

[BB24] Alessandro Barbarino and Efstathia Bura. Forecasting nearequivalence of linear dimension reduction methods in large panels of macro-variables. Econometrics and Statistics, 31 (??):1–18, July 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001222.

Bartolucci:2017:MTR

[BBP17] Francesco Bartolucci, Silvia Bacci, and Claudia Pigini. Misspecification test for random effects in generalized linear finite-mixture models for clustered binary and ordered data. Econometrics and Statistics, 3(??):112–131, July 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306216300314.

Bauwens:2017:DCM

[BBS17] Luc Bauwens, Manuela Braione, and Giuseppe Storti. A dynamic component model for forecasting high-dimensional realized covariance matrices. Econometrics and Statistics, 1 (??):40–61, January 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300132.

Billio:2024:CSF

[BCCI24] Monica Billio, Roberto Casarin, Michele Costola, and Matteo Iacopini. COVID-19 spreading in financial networks: a semiparametric matrix regression model. Econometrics and Statistics, 29 (??):113–131, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001180.

Billio:2023:NRS

[BCFP23] Monica Billio, Massimiliano Caporin, Lorenzo Frattarolo, and Loriana Pelizzon. Networks in risk spillovers: a multivariate GARCH perspective. Econometrics and Statistics, 28 (??):1–29, October 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000058.

Bravo:2017:GEL

[BCJC17] Francesco Bravo, Ba M. Chu, and David T. Jacho-Chávez. Generalized empirical likelihood m testing for semiparametric models with time series data. Econometrics and Statistics, 4 (??):18–30, October 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300023.

Baillie:2024:CLS

[BCR24] Richard T. Baillie, Dooyeon Cho, and Seunghwa Rho. Combining long and short memory in time series models: the role of asymptotic correlations of the MLEs. Econometrics and Statistics, 29(??):88–112, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000752.

Bleher:2022:KMA

[BD22] Johannes Bleher and Thomas Dimpfl. Knitting multi-annual high-frequency Google trends to predict inflation and consumption. Econometrics and Statistics, 24(??):1–26, October 2022. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306221001210.

Bergsma:2020:RP

[Ber20] Wicher P. Bergsma. Regression with I-priors. Econometrics and Statistics, 14(??):89–111, April 2020. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306219300632.

Beutner:2023:REA

[Beu23] Eric Beutner. A review of effective age models and associated non- and semiparametric methods. Econometrics and Statistics, 28(??):105–119, October 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001490.

Bach:2018:SCD

[BFS18] Philipp Bach, Helmut Farbmacher, and Martin Spindler. Semiparametric count data modeling with an application to health service demand. Econometrics and Statistics, 8(??):

125–140, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300710.

Boudt:2024:RIF

[BH24] Kris Boudt and Ewoud Heyndels. Robust interactive fixed effects. Econometrics and Statistics, 29(??):206–223, January 2024. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306222000028.

Burdejova:2017:CPT

[BHKX17] P. Burdejova, W. Härdle, P. Kokoszka, and Q. Xiong. Change point and trend analyses of annual expectile curves of tropical storms. Econometrics and Statistics, 1(??):101–117, January 2017. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306216300120.

Boswijk:2017:SIT

[BHL^{+}17] Peter Boswijk, Marc Hallin, Degui Li, Dimitris N. Politis, and Robert Taylor. Special issue on time series econometrics. Econometrics and Statistics, 4(??):1–2, October 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300412.

Brown:2019:STD

[BI19] Donald Brown and Rustam Ibragimov. Sign tests for dependent observations. Econometrics and Statistics, 10(??):1–8, April 2019. CODEN ???? ISSN 2452-3062. URL https://www.

sciencedirect.com/science/article/pii/S2452306218300935.

Bindele:2018:CMR

[Bin18] Huybrechts F. Bindele. Covariates missing at random under signed-rank inference. Econometrics and Statistics, 8(??): 78–93, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300315.

Becker:2018:FRC

[BK18] Martin Becker and Stefan Klößner. Fast and reliable computation of generalized synthetic controls. Econometrics and Statistics, 5(??):1–19, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300692.

Breitung:2022:BCM

[BKH22] Jörg Breitung, Sebastian Kripfganz, and Kazuhiko Hayakawa. Bias-corrected method of moments estimators for dynamic panel data models. Econometrics and Statistics, 24(??): 116–132, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000770.

Broda:2018:AES

[BKP18] Simon A. Broda, Jochen Krause, and Marc S. Paolella. Approximating expected shortfall for heavy-tailed distributions. Econometrics and Statistics, 8(??):184–203, October 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300643.

Blackburn:2022:TCD

[Bla22] McKinley L. Blackburn. Testing for coefficient differences across nested linear regression specifications. Econometrics and Statistics, 23(??):1–18, July 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000496.

Blasques:2021:FSO

[BLvV21] Francisco Blasques, André Lucas, and Andries C. van Vlodrop. Finite sample optimality of score-driven volatility models: Some Monte Carlo evidence. Econometrics and Statistics, 19(??):47–57, July 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300435.

Bladt:2022:TSC

[BM22] Martin Bladt and Alexander J. McNeil. Time series copula models using d-vines and v-transforms. Econometrics and Statistics, 24(??):27–48, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000800.

Berger:2022:SPE

[BP22] Yves G. Berger and Valentin Patilea. A semi-parametric empirical likelihood approach for conditional estimating equations under endogenous selection. Econometrics and Statistics, 24 (??):151–163, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001489.

Breitung:2018:ACD

[BS18] Jörg Breitung and Sven Schreiber. Assessing causality and delay within a frequency band. Econometrics and Statistics, 6(??):57–73, April 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300370.

Beare:2019:IBT

[BS19] Brendan K. Beare and Xiaoxia Shi. An improved bootstrap test of density ratio ordering. Econometrics and Statistics, 10(??):9–26, April 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300510.

Behrendt:2021:NAG

[BS21] Simon Behrendt and Karsten Schweikert. A note on adaptive group lasso for structural break time series. Econometrics and Statistics, 17(??):156–172, January 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300320.

Borrotti:2023:MOO

[BSM23] Matteo Borrotti, Francesco Sambo, and Kalliopi Mylona. Multiobjective optimisation of split-plot designs. Econometrics and Statistics, 28(??):163–172, October 2023. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306222000417.

Battagliola:2022:BAE

[BSTS22] Maria Laura Battagliola, Helle Sørensen, Anders Tolver, and Ana-Maria Staicu. A bias-adjusted estimator in quantile regression for clustered data. Econometrics and Statistics, 23 (??):165–186, July 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000794.

Bernard:2024:SMS

[BV24] Gaspard Bernard and Thomas Verdebout. On some multivariate sign tests for scatter matrix eigenvalues. Econometrics and Statistics, 29(??):252–260, January 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000472.

Catani:2017:CLM

[CA17] P. S. Catani and N. J. C. Ahlgren. Combined Lagrange multiplier test for ARCH in vector autoregressive models. Econometrics and Statistics, 1(??):62–84, January 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300107.

Chen:2023:NMF

[CCCW23] Ping-Yang Chen, Ray-Bing Chen, Yu-Shi Chen, and Weng Kee Wong. Numerical methods for finding A-optimal designs analytically. Econometrics and Statistics, 28(??):155–162, October 2023. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306222000867.

Chang:2024:EBM

[CCW24] Xiao-Wen Chang, Zhilong Chen, and Jinming Wen. An extended Babai method for estimating linear model based integer parameters. Econometrics and Statistics, 29(??):238– 251, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000089.

Castle:2023:RDR

[CDH23] Jennifer L. Castle, Jurgen A. Doornik, and David F. Hendry. Robust discovery of regression models. Econometrics and Statistics, 26(??):31–51, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000629.

Cantoni:2020:SIM

[CdL20] Eva Cantoni and Xavier de Luna. Semiparametric inference with missing data: Robustness to outliers and model misspecification. Econometrics and Statistics, 16(??):108–120, October 2020. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306220300198.

Centorrino:2021:NIV

[CF21] Samuele Centorrino and Jean-Pierre Florens. Nonparametric instrumental variable estimation of binary response models with continuous endogenous regressors. Econometrics and Statistics, 17(??):35–63, January 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622030071X.

Cipollini:2019:MES

[CG19] Fabrizio Cipollini and Giampiero M. Gallo. Modeling Euro STOXX 50 volatility with common and market-specific components. Econometrics and Statistics, 11(??):22–42, July 2019. CO-

DEN ???? ISSN 2452-3062. URL https://www.sciencedirect.
com/science/article/pii/S245230621830073X.

Chan:2022:TAS

[Cha22] Wai-Sum Chan. On temporal aggregation of some nonlinear time-series models. Econometrics and Statistics, 21(??): 38–49, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300411.

Chaurasia:2023:CRF

[Cha23] Ashok Chaurasia. Combining rules for F - and beta-statistics from multiply-imputed data. Econometrics and Statistics, 25 (??):51–65, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001076.

Cheng:2024:EES

[CHM24] Yuzhong Cheng, Nicole Hufnagel, and Hiroki Masuda. Estimation of ergodic square-root diffusion under high-frequency sampling. Econometrics and Statistics, 32(??):73–87, October 2024. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306222000727.

Czado:2019:MTD

[CIO19] Claudia Czado, Eugen Ivanov, and Yarema Okhrin. Modelling temporal dependence of realized variances with vines. Econometrics and Statistics, 12(??):198–216, October 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306219300218.

Cantoni:2024:RCM

[CJG24] Eva Cantoni, Nadège Jacot, and Paolo Ghisletta. Review and comparison of measures of explained variation and model selection in linear mixed-effects models. Econometrics and Statistics, 29(??):??, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000630.

Cizek:2021:JPV

[CK21a] Pavel Cízek and Chao Hui Koo. Jump-preserving varyingcoefficient models for nonlinear time series. Econometrics and Statistics, 19(??):58–96, July 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306220300472.

Colubi:2021:AESa

[CK21b] Ana Colubi and Erricos Kontoghiorghes. Advances of Econometrics and Statistics (EcoSta), 1st issue. Econometrics and Statistics, 18(??):28, April 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000356.

Colubi:2021:AESb

[CK21c] Ana Colubi and Erricos Kontoghiorghes. Advances of Econometrics and Statistics (EcoSta), 2nd issue. Econometrics and Statistics, 20(??):1, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000927.

Cho:2022:HDG

[CK22] Haeran Cho and Karolos K. Korkas. High-dimensional GARCH process segmentation with an application to value-at-risk. Econometrics and Statistics, 23(??):187–203, July 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230622100085X.

Colubi:2023:ESI

[CK23] Ana Colubi and Erricos Kontoghiorghes. Editorial special issues on the 20th anniversary of the CMStatistics (Computational and Methodological Statistics). Econometrics and Statistics, 26(??):1–2, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306223000126.

Cho:2024:DSA

[CK24] Haeran Cho and Claudia Kirch. Data segmentation algorithms: Univariate mean change and beyond. Econometrics and Statistics, 30(??):76–95, April 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001234.

Craig:2023:CPR

[CKL^{+}23] Sarah J. C. Craig, Ana M. Kenney, Junli Lin, Ian M. Paul, Leann L. Birch, Jennifer S. Savage, Michele E. Marini, Francesca Chiaromonte, Matthew L. Reimherr, and Kateryna D. Makova. Constructing a polygenic risk score for childhood obesity using functional data analysis. Econometrics and Statistics, 25 (??):66–86, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001295.

Chronopoulos:2021:KBV

[CKP21] Ilias Chronopoulos, George Kapetanios, and Katerina Petrova. Kernel-based volatility generalised least squares. Econometrics and Statistics, 20(??):2–11, October 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306219300644.

Canale:2023:ISS

[CLNP23] Antonio Canale, Antonio Lijoi, Bernardo Nipoti, and Igor Prünster. Inner spike and slab Bayesian nonparametric models. Econometrics and Statistics, 27(??):120–135, July 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230622100143X.

Curato:2018:SVE

[CMR18] Imma Valentina Curato, Maria Elvira Mancino, and Maria Cristina Recchioni. Spot volatility estimation using the Laplace transform. Econometrics and Statistics, 6(??):22–43, April 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230621630017X.

Corradin:2022:OSS

[CNBN22] Riccardo Corradin, Luis Enrique Nieto-Barajas, and Bernardo Nipoti. Optimal stratification of survival data via Bayesian nonparametric mixtures. Econometrics and Statistics, 22 (??):17–38, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000605.

Colombi:2020:STP

[Col20] Roberto Colombi. Selection tests for possibly misspecified hierarchical multinomial marginal models. Econometrics and Statistics, 16(??):136–147, October 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300449.

Choi:2017:SIB

[COSW17] Taeryon Choi, Yasuhiro Omori, Michael Smith, and Stephen G. Walker. Special issue on Bayesian methods in statistics and econometrics. Econometrics and Statistics, 3(??):1–2, July 2017. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306217300400.

Chu:2023:MRR

[COySS23] Amanda M. Y. Chu, Yasuhiro Omori, Hing yu So, and Mike K. P. So. A multivariate randomized response model for sensitive binary data. Econometrics and Statistics, 27(??):16–35, July 2023. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306222000041.

Characiejus:2020:GWN

[CR20] Vaidotas Characiejus and Gregory Rice. A general white noise test based on kernel lag-window estimates of the spectral density operator. Econometrics and Statistics, 13(??): 175–196, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300073.

Creel:2017:NNI

[Cre17] Michael Creel. Neural nets for indirect inference. Econometrics and Statistics, 2(??):36–49, April 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306216300326.

Colubi:2023:FSF

[CRG23] Ana Colubi and Ana Belén Ramos-Guajardo. Fuzzy sets and (fuzzy) random sets in Econometrics and Statistics. Econometrics and Statistics, 26(??):84–98, April 2023. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306222000788.

Castagnetti:2019:TSE

[CRT19] Carolina Castagnetti, Eduardo Rossi, and Lorenzo Trapani. A two-stage estimator for heterogeneous panel models with common factors. Econometrics and Statistics, 11(??):63–82, July 2019. CODEN ???? ISSN 2452-3062. URL https://www.

sciencedirect.com/science/article/pii/S245230621730093X.

Curato:2022:SLE

[CS22] Imma Valentina Curato and Simona Sanfelici. Stochastic leverage effect in high-frequency data: a Fourier based analysis. Econometrics and Statistics, 23(??):53–82, July 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000319.

Chen:2023:NSB

[CW23] Su Chen and Stephen G. Walker. A new statistic for Bayesian hypothesis testing. Econometrics and Statistics, 26 (??):139–152, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001246.

Chen:2023:BER

[CWL23] Cathy W. S. Chen, Toshiaki Watanabe, and Edward M. H. Lin. Bayesian estimation of realized GARCH-type models with application to financial tail risk management. Econometrics and Statistics, 28(??):30–46, October 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000484.

Cong:2021:LRT

[CY21] Lin Cong and Weixin Yao. A likelihood ratio test of a homoscedastic multivariate normal mixture against a heteroscedastic multivariate normal mixture. Econometrics and Statistics, 18(??):79–88, April 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000046.

Chai:2017:IGE

[CZJ^{+}17] Hao Chai, Qingzhao Zhang, Yu Jiang, Guohua Wang, Sanguo Zhang, Syed Ejaz Ahmed, and Shuangge Ma. Identifying geneenvironment interactions for prognosis using a robust approach. Econometrics and Statistics, 4(??):105–120, October 2017. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306216300065.

Czudaj:2019:DBT

[Czu19] Robert L. Czudaj. Dynamics between trading volume, volatility and open interest in agricultural futures markets: a Bayesian time-varying coefficient approach. Econometrics and Statistics, 12(??):78–145, October 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306219300267.

Castro:2021:ASL

[dBCR21] Tomás del Barrio Castro and Heiko Rachinger. Aggregation of seasonal long-memory processes. Econometrics and Statistics, 17(??):95–106, January 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300691.

Dirick:2022:HMC

[DCVB22] Lore Dirick, Gerda Claeskens, Andrey Vasnev, and Bart Baesens. A hierarchical mixture cure model with unobserved heterogeneity for credit risk. Econometrics and Statistics, 22 (??):39–55, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000022.

DeBlander:2020:IEC

[De 20] Rembert De Blander. Iterative estimation correcting for error auto-correlation in short panels, applied to lagged dependent variable models. Econometrics and Statistics, 15 (??):3–29, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300186.

Demetriou:2021:ADB

[Dem21] I. C. Demetriou. A O(n) algorithm for the discrete best L_{4} monotonic approximation problem. Econometrics and Statistics, 17 (??):130–144, January 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300393.

Dimitriou-Fakalou:2019:AEE

[DF19] Chrysoula Dimitriou-Fakalou. On accepting the edge-effect (for the inference of ARMA-type processes in Z^{2}). Econometrics and Statistics, 10(??):53–70, April 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300091.

DiIorio:2021:ERC

[DF21] Francesca Di Iorio and Stefano Fachin. Evaluating restricted common factor models for non-stationary data. Econometrics and Statistics, 17(??):64–75, January 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306220300873.

Daouia:2020:RFE

[DFS20] Abdelaati Daouia, Jean-Pierre Florens, and Léopold Simar. Robust frontier estimation from noisy data: a Tikhonov regularization approach. Econometrics and Statistics, 14 (??):1–23, April 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300492.

Dette:2022:CIP

[DGK22] Holger Dette, Vasyl Golosnoy, and Janosch Kellermann. Correcting intraday periodicity bias in realized volatility measures. Econometrics and Statistics, 23(??):36–52, July 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000320.

Dufour:2020:ESI

[DHW20] Jean-Marie Dufour, Alain Hecq, and Alan Wan. EcoSta special issue on theoretical econometrics. Econometrics and Statistics, 15(??):1–2, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300459.

Dai:2020:BLS

[DJF20] Ning Dai, Galin L. Jones, and Mark Fiecas. Bayesian longitudinal spectral estimation with application to restingstate fMRI data analysis. Econometrics and Statistics, 15 (??):104–116, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300061.

Deistler:2017:NIV

[DKA17] Manfred Deistler, Lukas Koelbl, and Brian D. O. Anderson. Non-identifiability of VMA and VARMA systems in the mixed frequency case. Econometrics and Statistics, 4 (??):31–38, October 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300302.

Dohler:2018:DMB

[Döh18] Sebastian Döhler. A discrete modification of the Benjamini– Yekutieli procedure. Econometrics and Statistics, 5(??):137– 147, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300351.

DiCiccio:2019:IWL

[DRW19] Cyrus J. DiCiccio, Joseph P. Romano, and Michael Wolf. Improving weighted least squares inference. Econometrics and Statistics, 10(??):96–119, April 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300364.

deValk:2018:HQE

[dVC18] Cees de Valk and Juan-Juan Cai. A high quantile estimator based on the log-generalized Weibull tail limit. Econometrics and Statistics, 6(??):107–128, April 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300114.

Bermudez:2024:INL

[dZBMRV24] P. de Zea Bermudez, J. Miguel Marín, Håvard Rue, and Helena Veiga. Integrated nested Laplace approximations for threshold stochastic volatility models. Econometrics and Statistics, 30(??):15–35, April 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001003.

Eguchi:2018:MCG

[Egu18] Shoichi Eguchi. Model comparison for generalized linear models with dependent observations. Econometrics and Statistics, 5 (??):171–188, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300357.

Emura:2023:CPE

[ELS23] Takeshi Emura, Ching-Chieh Lai, and Li-Hsien Sun. Change point estimation under a copula-based Markov chain model for binomial time series. Econometrics and Statistics, 28(??): 120–137, October 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000836.

ElMethni:2018:IEE

[ES18] Jonathan El Methni and Gilles Stupfler. Improved estimators of extreme Wang distortion risk measures for very heavy-tailed distributions. Econometrics and Statistics, 6(??): 129–148, April 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300151.

Ferraro:2024:FMH

[Fer24] Maria Brigida Ferraro. Fuzzy k-means: history and applications. Econometrics and Statistics, 30(??):110–123, April 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001398.

Fontaine:2020:MNL

[FFO20] Charles Fontaine, Ron D. Frostig, and Hernando Ombao. Modeling non-linear spectral domain dependence using copulas with applications to rat local field potentials. Econometrics and Statistics, 15(??):85–103, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300450.

Ferraro:2021:CTM

[FGV21] Maria Brigida Ferraro, Paolo Giordani, and Maurizio Vichi. A class of two-mode clustering algorithms in a fuzzy setting. Econometrics and Statistics, 18(??):63–78, April 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300381.

Funke:2019:NET

[FH19] Benedikt Funke and Masayuki Hirukawa. Nonparametric estimation and testing on discontinuity of positive supported densities: a kernel truncation approach. Econometrics and Statistics, 9(??):156–170, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300679.

Funke:2021:BCL

[FH21] Benedikt Funke and Masayuki Hirukawa. Bias correction for local linear regression estimation using asymmetric kernels via the skewing method. Econometrics and Statistics, 20(??):

109–130, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300204.

Filippeli:2020:DBP

[FHT20] Thomai Filippeli, Richard Harrison, and Konstantinos Theodoridis. DSGE-based priors for BVARs and quasi-Bayesian DSGE estimation. Econometrics and Statistics, 16(??):1–27, October 2020. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306219300012.

Ferraty:2022:SIF

[FKKP22] Frederic Ferraty, Alois Kneip, Piotr Kokoszka, and Alexander Petersen. 2nd special issue on functional data analysis. Econometrics and Statistics, 21(??):112–113, January 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001350.

Fazzi:2024:BCV

[FKM24] Antonio Fazzi, Alexander Kukush, and Ivan Markovsky. Bias correction for Vandermonde low-rank approximation. Econometrics and Statistics, 31(??):38–48, July 2024. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001088.

Florios:2018:HIS

[Flo18] Kostas Florios. A hyperplanes intersection simulated annealing algorithm for maximum score estimation. Econometrics and Statistics, 8(??):37–55, October 2018. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306217300291.

Fokianos:2017:BTS

[FM17] Konstantinos Fokianos and Theodoros Moysiadis. Binary time series models driven by a latent process. Econometrics and Statistics, 2(??):117–130, April 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300096.

Figini:2020:MRI

[FMU20] Silvia Figini, Mario Maggi, and Pierpaolo Uberti. The market rank indicator to detect financial distress. Econometrics and Statistics, 14(??):63–73, April 2020. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306218300017.

Fokianos:2024:MCT

[Fok24] Konstantinos Fokianos. Multivariate count time series modelling. Econometrics and Statistics, 31(??):100–116, July 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001374.

Forcina:2017:FSA

[For17] Antonio Forcina. A Fisher-scoring algorithm for fitting latent class models with individual covariates. Econometrics and Statistics, 3(??):132–140, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300089.

Fan:2017:HDA

[FR17] Zhaohu Fan and Matthew Reimherr. High-dimensional adaptive function-on-scalar regression. Econometrics and Statistics, 1(??):167–183, January 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300053.

Fitzpatrick:2022:AMC

[FS22] Matthew Fitzpatrick and Michael Stewart. Asymptotics for Markov chain mixture detection. Econometrics and Statistics, 22(??):56–66, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001337.

Follett:2019:APB

[FY19] Lendie Follett and Cindy Yu. Achieving parsimony in Bayesian vector autoregressions with the horseshoe prior. Econometrics and Statistics, 11(??):130–144, July 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306219300036.

Ferraty:2019:NRC

[FZF19] Frédéric Ferraty, Anthony Zullo, and Mathieu Fauvel. Nonparametric regression on contaminated functional predictor with application to hyperspectral data. Econometrics and Statistics, 9 (??):95–107, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300138.

Geenens:2022:NCA

[GD22] Gery Geenens and Richard Dunn. A nonparametric copula approach to conditional value-at-risk. Econometrics and Statistics, 21(??):19–37, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300666.

Garcia-Enriquez:2019:LWE

[GEH19] Javier García-Enríquez and Javier Hualde. Local Whittle estimation of long memory: Standard versus bias-reducing techniques. Econometrics and Statistics, 12(??):66–77, October 2019. CO-

DEN ???? ISSN 2452-3062. URL https://www.sciencedirect.
com/science/article/pii/S2452306219300280.

Grillenzoni:2019:LPD

[GF19] Carlo Grillenzoni and Michele Fornaciari. On-line peak detection in medical time series with adaptive regression methods. Econometrics and Statistics, 10(??):134–150, April 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306218300480.

Gattone:2022:SCB

[GFED22] Stefano Antonio Gattone, Francesca Fortuna, Adelia Evangelista, and Tonio Di Battista. Simultaneous confidence bands for the functional mean of convex curves. Econometrics and Statistics, 24(??):183–193, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001441.

Grossmann:2023:POB

[GG23] Heiko Großmann and Steven G. Gilmour. Partially orthogonal blocked three-level response surface designs. Econometrics and Statistics, 28(??):138–154, October 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001015.

Goegebeur:2023:WTE

[GGHQ23] Yuri Goegebeur, Armelle Guillou, Nguyen Khanh Le Ho, and Jing Qin. A Weissman-type estimator of the conditional marginal expected shortfall. Econometrics and Statistics, 27 (??):173–196, July 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001131.

Gorecki:2018:CPD

[GHK18] Tomasz Górecki, Lajos Horváth, and Piotr Kokoszka. Change point detection in heteroscedastic time series. Econometrics and Statistics, 7(??):63–88, July 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300667.

Gourieroux:2019:RAM

[GJ19] Christian Gourieroux and Joann Jasiak. Robust analysis of the martingale hypothesis. Econometrics and Statistics, 9 (??):17–41, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300479.

Genest:2019:IST

[GKD19] Christian Genest, Ivan Kojadinovic, and Fabrizio Durante. Introduction to the special topic on copula modeling. Econometrics and Statistics, 12(??):146–147, October 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306219300462.

Gissibl:2018:TDR

[GKO18] Nadine Gissibl, Claudia Klüppelberg, and Moritz Otto. Tail dependence of recursive max-linear models with regularly varying noise variables. Econometrics and Statistics, 6(??): 149–167, April 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621830008X.

Gourieroux:2018:CII

[GM18] C. Gourieroux and A. Monfort. Composite indirect inference with application to corporate risks. Econometrics and Statistics, 7(??):30–45, July 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300886.

Gourieroux:2021:MRM

[GM21] C. Gourieroux and A. Monfort. Model risk management: Valuation and governance of pseudo-models. Econometrics and Statistics, 17(??):1–22, January 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300708.

Ghosal:2022:SBT

[GM22] Rahul Ghosal and Arnab Maity. A score based test for functional linear concurrent regression. Econometrics and Statistics, 21(??):114–130, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000617.

Grigoryeva:2018:VFU

[GOP18] Lyudmila Grigoryeva, Juan-Pablo Ortega, and Anatoly Peresetsky. Volatility forecasting using global stochastic financial trends extracted from non-synchronous data. Econometrics and Statistics, 5(??):67–82, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300059.

Guisinger:2024:ICB

[GOS24] Amy Y. Guisinger, Michael T. Owyang, and Daniel Soques. Industrial connectedness and business cycle comovements. Econometrics and Statistics, 29(??):132–149, January 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000988.

Gambacciani:2017:RNM

[GP17] Marco Gambacciani and Marc S. Paolella. Robust normal mixtures for financial portfolio allocation. Econometrics and Statistics, 3(??):91–111, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300126.

Galharret:2023:BAM

[GP23] Jean-Michel Galharret and Anne Philippe. Bayesian analysis for mediation and moderation using g-priors. Econometrics and Statistics, 27(??):161–172, July 2023. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001635.

Ghysels:2019:EMR

[GQ19] Eric Ghysels and Hang Qian. Estimating MIDAS regressions via OLS with polynomial parameter profiling. Econometrics and Statistics, 9(??):1–16, January 2019. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306218300066.

Guindani:2023:ESI

[GR23] Michele Guindani and F. Javier Rubio. Editorial: Special issue on biostatistics. Econometrics and Statistics, 25 (??):49–50, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622200096X.

Gonzalez-Rodriguez:2017:CBM

[GRC17] Gil González-Rodríguez and Ana Colubi. On the consistency of bootstrap methods in separable Hilbert spaces. Econometrics and Statistics, 1(??):118–127, January 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300259.

Golosnoy:2020:SIR

[GSSL20] Vasyl Golosnoy, Wolfgang Schmid, Miriam Isabel Seifert, and Taras Lazariv. Statistical inferences for realized portfolio weights. Econometrics and Statistics, 14(??):49–62, April 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306218300534.

Girard:2022:FEE

[GSUC22] Stéphane Girard, Gilles Stupfler, and Antoine Usseglio-Carleve. Functional estimation of extreme conditional expectiles. Econometrics and Statistics, 21(??):131–158, January 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000642.

Gao:2023:TWT

[GT23] Zhaoxing Gao and Ruey S. Tsay. A two-way transformed factor model for matrix-variate time series. Econometrics and Statistics, 27(??):83–101, July 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001027.

Guillou:2018:SIS

[Gui18] Armelle Guillou. Special issue on statistics of extremes and applications. Econometrics and Statistics, 6(??):106, April 2018. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306218300108.

Gruber:2017:BOV

[GW17] Lutz F. Gruber and Mike West. Bayesian online variable selection and scalable multivariate volatility forecasting in simultaneous graphical dynamic linear models. Econometrics and Statistics, 3(??):3–22, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300163.

Hauzenberger:2021:FMP

[Hau21] Niko Hauzenberger. Flexible mixture priors for large timevarying parameter models. Econometrics and Statistics, 20 (??):87–108, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000654.

Hayakawa:2019:AIR

[Hay19] Kazuhiko Hayakawa. Alternative over-identifying restriction test in the GMM estimation of panel data models. Econometrics and Statistics, 10(??):71–95, April 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300297.

Hayakawa:2024:RDC

[Hay24] Kazuhiko Hayakawa. Recent development of covariance structure analysis in economics. Econometrics and Statistics, 29 (??):??, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001179.

Hu:2020:HBM

[HGFO20] Lechuan Hu, Michele Guindani, Norbert J. Fortin, and Hernando Ombao. A hierarchical Bayesian model for differential connectivity in multi-trial brain signals. Econometrics and Statistics, 15(??):117–135, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300423.

Hinde:2017:SIM

[HILM17] John Hinde, Salvatore Ingrassia, Tsung-I Lin, and Paul D. Mc-Nicholas. Special issue on mixture models. Econometrics and Statistics, 3(??):89–90, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300394.

Hirukawa:2023:RCM

[Hir23] Masayuki Hirukawa. Robust covariance matrix estimation in time series: a review. Econometrics and Statistics, 27 (??):36–61, July 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001428.

Hadjiantoni:2022:ANM

[HK22] Stella Hadjiantoni and Erricos John Kontoghiorghes. An alternative numerical method for estimating large-scale time-varying parameter seemingly unrelated regressions models. Econometrics and Statistics, 21(??):1–18, January 2022. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221000095.

He:2018:SIQ

[HKLW18] Xuming He, Thomas Kneib, Carlos Lamarche, and Lan Wang. Special issue on quantile regression and semiparametric methods. Econometrics and Statistics, 8(??):1–2, October 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306218300558.

He:2019:SSM

[HKTZ19] Changli He, Jian Kang, Timo Teräsvirta, and Shuhua Zhang. The shifting seasonal mean autoregressive model and seasonality in the Central England monthly temperature series, 1772–2016. Econometrics and Statistics, 12(??):1–24, October 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306219300292.

Harvey:2023:DTM

[HL23] Andew Harvey and Yin Liao. Dynamic tobit models. Econometrics and Statistics, 26(??):72–83, April 2023. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001064.

Huang:2020:CES

[HLU20] Bai Huang, Tae-Hwy Lee, and Aman Ullah. Combined estimation of semiparametric panel data models. Econometrics and Statistics, 15(??):30–45, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300255.

Hofert:2018:VDH

[HO18]  Marius Hofert and Wayne Oldford.  Visualizing depen-

dence in high-dimensional data: An application to S&P 500 constituent data. Econometrics and Statistics, 8(??): 161–183, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621730031X.

Hristache:2021:EMO

[HP21] Marian Hristache and Valentin Patilea. Equivalent models for observables under the assumption of missing at random. Econometrics and Statistics, 20(??):153–165, October 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300332.

Hans:2023:EBM

[HPW23] Christopher M. Hans, Mario Peruggia, and Junyan Wang. Empirical Bayes model averaging with influential observations: Tuning Zellner’s g prior for predictive robustness. Econometrics and Statistics, 27(??):102–119, July 2023. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001477.

Hofert:2022:MTS

[HPZ22] Marius Hofert, Avinash Prasad, and Mu Zhu. Multivariate time-series modeling with generative neural networks. Econometrics and Statistics, 23(??):147–164, July 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230622100126X.

Hlouskova:2020:GEA

[HS20] Jaroslava Hlouskova and Leopold Sögner. GMM estimation of affine term structure models. Econometrics and Statistics, 13 (??):2–15, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300620.

Haupt:2018:EGT

[HSS18] Harry Haupt, Joachim Schnurbus, and Willi Semmler. Estimation of grouped, time-varying convergence in economic growth. Econometrics and Statistics, 8(??):141–158, October 2018. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306217300722.

Hallin:2023:FVR

[HT23] Marc Hallin and Carlos Trucíos. Forecasting value-at-risk and expected shortfall in large portfolios: a general dynamic factor model approach. Econometrics and Statistics, 27(??):1–15, July 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000563.

Hu:2021:SVS

[Hu21] Guanyu Hu. Spatially varying sparsity in dynamic regression models. Econometrics and Statistics, 17(??):23–34, January 2021. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306220300861.

Hecq:2021:FBM

[HV21] Alain Hecq and Elisa Voisin. Forecasting bubbles with mixed causal-noncausal autoregressive models. Econometrics and Statistics, 20(??):29–45, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622030040X.

Hasnat:2017:ECC

[HVBJ17] Md. Abul Hasnat, Julien Velcin, Stephane Bonnevay, and Julien Jacques. Evolutionary clustering for categorical data using parametric links among multinomial mixture models. Econometrics and Statistics, 3(??):141–159, July 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306217300175.

Hafner:2024:DPS

[HW24] Christian M. Hafner and Linqi Wang. Dynamic portfolio selection with sector-specific regularization. Econometrics and Statistics, 32(??):17–33, October 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000016.

Ingrassia:2022:SIM

[IL22] Salvatore Ingrassia and Tsung-I Lin. The 2nd special issue on mixture models. Econometrics and Statistics, 22 (??):1–2, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000235.

Jiang:2017:LDQ

[Jia17] Wenxin Jiang. On limiting distribution of quasi-posteriors under partial identification. Econometrics and Statistics, 3 (??):60–72, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300308.

Javed:2024:EEM

[JLM24] Farrukh Javed, Nicola Loperfido, and Stepan Mazur. Edgeworth expansions for multivariate random sums. Econometrics and Statistics, 31(??):66–80, July 2024. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221000551.

Jaskowski:2021:SCS

[JM21] Marcin Jaskowski and Michael McAleer. Spurious cross-sectional dependence in credit spread changes. Econometrics and Statistics, 18(??):12–27, April 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300619.

Johnson:2020:SIN

[JS20]  Timothy D. Johnson and Armin Schwartzman.  Special

issue on neuroimaging. Econometrics and Statistics, 15 (??):84, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300447.

Kwak:2024:DPG

[KALP24] Seung Woo Kwak, Jeongyoun Ahn, Jaewoo Lee, and Cheolwoo Park. Differentially private goodness-of-fit tests for continuous variables. Econometrics and Statistics, 31(??):81–99, July 2024. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306221001143.

Karame:2018:NPF

[Kar18] Frédéric Karamé. A new particle filtering approach to estimate stochastic volatility models with Markov-switching. Econometrics and Statistics, 8(??):204–230, October 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306218300352.

Kawka:2022:CSD

[Kaw22] Rafael Kawka. Convergence of spectral density estimators in the locally stationary framework. Econometrics and Statistics, 24(??):94–115, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300678.

Kreuzer:2021:BIS

[KC21] Alexander Kreuzer and Claudia Czado. Bayesian inference for a single factor copula stochastic volatility model using Hamiltonian Monte Carlo. Econometrics and Statistics, 19 (??):130–150, July 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000010.

Kartsonaki:2023:RRR

[KC23] Christiana Kartsonaki and D. R. Cox. Regression reconstruction from a retrospective sample. Econometrics and Statistics, 25(??):87–92, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622030085X.

Ko:2019:CIC

[KH19] Vinnie Ko and Nils Lid Hjort. Copula information criterion for model selection with two-stage maximum likelihood estimation. Econometrics and Statistics, 12(??):167–180, October 2019. CO-

DEN ???? ISSN 2452-3062. URL https://www.sciencedirect.
com/science/article/pii/S245230621930005X.

Kiriliouk:2020:HTT

[Kir20] Anna Kiriliouk. Hypothesis testing for tail dependence parameters on the boundary of the parameter space. Econometrics and Statistics, 16(??):121–135, October 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306219300425.

Kiviet:2020:MDP

[Kiv20] Jan F. Kiviet. Microeconometric dynamic panel data methods: Model specification and selection issues. Econometrics and Statistics, 13(??):16–45, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300498.

Kiviet:2023:IFI

[Kiv23] Jan F. Kiviet. Instrument-free inference under confined regressor endogeneity and mild regularity. Econometrics and Statistics, 25(??):1–22, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001623.

Krupskii:2020:FCM

[KJ20] Pavel Krupskii and Harry Joe. Flexible copula models with dynamic dependence and application to financial data. Econometrics and Statistics, 16(??):148–167, October 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300216.

Klepsch:2017:PFA

[KKW17] J. Klepsch, C. Klüppelberg, and T. Wei. Prediction of functional ARMA processes with an application to traffic data. Econometrics and Statistics, 1(??):128–149, January 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230621630020X.

Kleppe:2022:ACF

[KLMO22] Tore Selland Kleppe, Roman Liesenfeld, Guilherme Valle Moura, and Atle Oglend. Analyzing commodity futures using factor state-space models with Wishart stochastic volatility. Econometrics and Statistics, 23(??):105–127, July 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000514.

Krishnamurthy:2018:FSV

[KLS18] Vikram Krishnamurthy, Elisabeth Leoff, and Jörn Sass. Filterbased stochastic volatility in continuous-time hidden Markov models. Econometrics and Statistics, 6(??):1–21, April 2018. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306216300144.

Kleppe:2017:ECS

[KO17] Tore Selland Kleppe and Atle Oglend. Estimating the competitive storage model: A simulated likelihood approach. Econometrics and Statistics, 4(??):39–56, October 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300333.

Kurose:2020:MBD

[KO20] Yuta Kurose and Yasuhiro Omori. Multiple-block dynamic equicorrelations with realized measures, leverage and endogeneity. Econometrics and Statistics, 13(??):46–68, January 2020. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306218300170.

Kokoszka:2017:SIF

[KOPS17] Piotr Kokoszka, Hanny Oja, Byeong Park, and Laura Sangalli. Special issue on functional data analysis. Econometrics and Statistics, 1(??):99–100, January 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306216300272.

Kiviet:2017:PTE

[KP17] Jan F. Kiviet and Milan Pleus. The performance of tests on endogeneity of subsets of explanatory variables scanned by simulation. Econometrics and Statistics, 2(??):1–21, April 2017. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306217300035.

Kapetanios:2021:SLW

[KPTV21] George Kapetanios, Simon Price, Menelaos Tasiou, and Alexia Ventouri. State-level wage Phillips curves. Econometrics and Statistics, 18(??):1–11, April 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S245230622030037X.

Kapetanios:2018:UFC

[KPY18] George Kapetanios, Simon Price, and Garry Young. A UK financial conditions index using targeted data reduction: Forecasting and structural identification. Econometrics and Statistics, 7(??):1–17, July 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300054.

Kasa:2022:IIG

[KR22] Siva Rajesh Kasa and Vaibhav Rajan. Improved inference of Gaussian mixture copula model for clustering and reproducibility analysis using automatic differentiation. Econometrics and Statistics, 22(??):67–97, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001040.

Kneib:2023:RAM

[KSS23] Thomas Kneib, Alexander Silbersdorff, and Benjamin Säfken. Rage against the mean — a review of distributional regression approaches. Econometrics and Statistics, 26(??): 99–123, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000824.

Koike:2019:OIS

[KT19] Yuta Koike and Yuta Tanoue. Oracle inequalities for sign constrained generalized linear models. Econometrics and Statistics, 11(??):145–157, July 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306219300085.

Kalogridis:2024:RPS

[KV24] Ioannis Kalogridis and Stefan Van Aelst. Robust penalized spline estimation with difference penalties. Econometrics and Statistics, 29(??):169–188, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000812.

Kontoghiorghes:2017:ES

[KVC17] Erricos Kontoghiorghes, Herman K. Van Dijk, and Ana Colubi. Econometrics and Statistics. Econometrics and Statistics, 1(??):1, January 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621630034X.

Lamperti:2018:ITC

[Lam18] Francesco Lamperti. An information theoretic criterion for empirical validation of simulation models. Econometrics and Statistics, 5(??):83–106, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300084.

Li:2022:SCE

[LBWM22] Hua Li, Zhidong Bai, Wing-Keung Wong, and Michael McAleer. Spectrally-corrected estimation for high-dimensional Markowitz mean-variance optimization. Econometrics and Statistics, 24 (??):133–150, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001209.

Li:2021:EBE

[LCD21] Dan Li, Adam Clements, and Christopher Drovandi. Efficient Bayesian estimation for GARCH-type models via sequential Monte Carlo. Econometrics and Statistics, 19 (??):22–46, July 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300319.

Lippi:2023:HDD

[LDA23] Marco Lippi, Manfred Deistler, and Brian Anderson. Highdimensional dynamic factor models: a selective survey and lines of future research. Econometrics and Statistics, 26 (??):3–16, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000302.

Li:2021:DCA

[LDL21] Linyuan Li, Pierre Duchesne, and Chu Pheuil Liou. On diagnostic checking in ARMA models with conditionally heteroscedastic martingale difference using wavelet methods. Econometrics and Statistics, 19(??):169–187, July 2021. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221000538.

Li:2017:NCI

[LEB17] Shu Li, Jan Ernest, and Peter Bühlmann. Nonparametric causal inference from observational time series through marginal integration. Econometrics and Statistics, 2(??):81–105, April 2017. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306216300260.

Liu-Evans:2018:UHO

[LEP18] Gareth Liu-Evans and Garry D. A. Phillips. On the use of higher order bias approximations for 2SLS and k-class estimators with non-normal disturbances and many instruments. Econometrics and Statistics, 6(??):90–105, April 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300448.

Ley:2023:WSF

[Ley23] Christophe Ley. When the score function is the identity function — a tale of characterizations of the normal distribution. Econometrics and Statistics, 26(??):153–160, April 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300836.

Li:2024:ELD

[Li24] Degui Li. Estimation of large dynamic covariance matrices: a selective review. Econometrics and Statistics, 29(??): 16–30, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000587.

Lutkepohl:2017:SVA

[LN17] Helmut Lütkepohl and Aleksei Netsunajev. Structural vector autoregressions with heteroskedasticity: A review of different volatility models. Econometrics and Statistics, 1(??): 2–18, January 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300223.

Leschinski:2019:MOS

[LS19] Christian Leschinski and Philipp Sibbertsen. Model order selection in periodic long memory models. Econometrics and Statistics, 9(??):78–94, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300042.

Lahiri:2020:FBM

[LS20] Ananya Lahiri and Rituparna Sen. Fractional Brownian markets with time-varying volatility and high-frequency data. Econometrics and Statistics, 16(??):91–107, October 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306218300765.

Liu:2021:BAH

[LS21] Hefei Liu and Xinyuan Song. Bayesian analysis of hidden Markov structural equation models with an unknown number of hidden states. Econometrics and Statistics, 18 (??):29–43, April 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300356.

Lutkepohl:2020:CJC

[LSBW20] Helmut Lütkepohl, Anna Staszewska-Bystrova, and Peter Winker. Constructing joint confidence bands for impulse response functions of VAR models — a review. Econometrics and Statistics, 13(??):69–83, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300741.

Lee:2024:FOS

[LSW24] JooChul Lee, Elizabeth D. Schifano, and HaiYing Wang. Fast optimal subsampling probability approximation for generalized linear models. Econometrics and Statistics, 29(??): 224–237, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000290.

Lamarche:2024:CQF

[LSY24] Carlos Lamarche, Xuan Shi, and Derek S. Young. Conditional quantile functions for zero-inflated longitudinal count data. Econometrics and Statistics, 31(??):49–65, July 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001106.

Lux:2022:INS

[Lux22] Thomas Lux. Inference for nonlinear state space models: a comparison of different methods applied to Markov-switching multifractal models. Econometrics and Statistics, 21(??): 69–95, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300307.

Liebl:2019:PRP

[LW19] Dominik Liebl and Fabian Walders. Parameter regimes in partial functional panel regression. Econometrics and Statistics, 11(??):105–115, July 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300339.

Leippold:2019:PFL

[LY19a] Markus Leippold and Hanlin Yang. Particle filtering, learning, and smoothing for mixed-frequency state-space models. Econometrics and Statistics, 12(??):25–41, October 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306219300437.

Li:2019:THH

[LY19b] Zhaoyuan Li and Jianfeng Yao. Testing for heteroscedasticity in high-dimensional regressions. Econometrics and Statistics, 9 (??):122–139, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300029.

Liu:2021:EIS

[LYBH21] Xin Liu, Grace Y. Yi, Glenn Bauman, and Wenqing He. Ensembling imbalanced-spatial-structured support vector machine. Econometrics and Statistics, 17(??):145–155, January 2021. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306220300344.

Maciak:2021:QLA

[Mac21a] Matús Maciak. Quantile LASSO in arbitrage-free option markets. Econometrics and Statistics, 18(??):106–116, April 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300526.

Maciak:2021:QLC

[Mac21b] Matús Maciak. Quantile LASSO with changepoints in panel data models applied to option pricing. Econometrics and Statistics, 20(??):166–175, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300046.

MacKinnon:2023:FCB

[Mac23] James G. MacKinnon. Fast cluster bootstrap methods for linear regression models. Econometrics and Statistics, 26 (??):52–71, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001404.

Marron:2017:BDC

[Mar17] J. S. Marron. Big Data in context and robustness against heterogeneity. Econometrics and Statistics, 2(??):73–80, April 2017. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306216300016.

Matsui:2020:QRF

[Mat20] Hidetoshi Matsui. Quadratic regression for functional response models. Econometrics and Statistics, 13(??):125– 136, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300024.

Maxand:2020:IIS

[Max20] Simone Maxand. Identification of independent structural shocks in the presence of multiple Gaussian components. Econometrics and Statistics, 16(??):55–68, October 2020. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306218300923.

Mayer:2022:LPS

[May22] Alexander Mayer. On the local power of some tests of strict exogeneity in linear fixed effects models. Econometrics and Statistics, 24(??):49–74, October 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000599.

Murray:2017:MSS

[MBM17] Paula M. Murray, Ryan P. Browne, and Paul D. McNicholas. A mixture of SDB skew-t factor analyzers. Econometrics and Statistics, 3(??):160–168, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300382.

Matilainen:2017:SDR

[MCNO17] M. Matilainen, C. Croux, K. Nordhausen, and H. Oja. Supervised dimension reduction for multivariate time series. Econometrics and Statistics, 4(??):57–69, October 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300345.

Mao:2020:ASV

[MCRV20] Xiuping Mao, Veronika Czellar, Esther Ruiz, and Helena Veiga. Asymmetric stochastic volatility models: Properties and particle filter-based simulated maximum likelihood estimation. Econometrics and Statistics, 13(??):84–105, January 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306219300486.

Melard:2022:IPA

[Mél22] Guy Mélard. An indirect proof for the asymptotic properties of VARMA model estimators. Econometrics and Statistics, 21 (??):96–111, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622100006X.

Muriel:2018:TND

[MGF18] Nelson Muriel and Graciela González-Farías. Testing the null of difference stationarity against the alternative of a stochastic unit root: A new test based on multivariate STUR. Econometrics and Statistics, 7(??):46–62, July 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300916.

Miller:2018:SRT

[Mil18] J. Isaac Miller. Simple robust tests for the specification of high-frequency predictors of a low-frequency series. Econometrics and Statistics, 5(??):45–66, January 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300119.

McCausland:2021:MSV

[MMP21] William McCausland, Shirley Miller, and Denis Pelletier. Multivariate stochastic volatility using the HESSIAN method. Econometrics and Statistics, 17(??):76–94, January 2021. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230622030068X.

Muschinski:2024:CBM

[MMS^{+}24] Thomas Muschinski, Georg J. Mayr, Thorsten Simon, Nikolaus Umlauf, and Achim Zeileis. Cholesky-based multivariate Gaussian regression. Econometrics and Statistics, 29(??): 261–281, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000168.

Morales-Onate:2021:BEL

[MOCB21] Víctor Morales-Oñate, Federico Crudu, and Moreno Bevilacqua. Blockwise Euclidean likelihood for spatio-temporal covariance models. Econometrics and Statistics, 20(??):176– 201, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000034.

Morana:2019:RSE

[Mor19] Claudio Morana. Regularized semiparametric estimation of high dimensional dynamic conditional covariance matrices. Econometrics and Statistics, 12(??):42–65, October 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306219300231.

Morana:2024:NMF

[Mor24] Claudio Morana. A new macro-financial condition index for the euro area. Econometrics and Statistics, 29(??): 64–87, January 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622100112X.

Marczak:2018:DCA

[MPG18] Martyna Marczak, Tommaso Proietti, and Stefano Grassi. A data-cleaning augmented Kalman filter for robust estimation of state space models. Econometrics and Statistics, 5(??): 107–123, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300102.

Maheu:2018:SIR

[MPSS18] John M. Maheu, Marc Paolella, Tak Kuen Siu, and Mike K. P. So. Special issue on risk management. Econometrics and Statistics, 8(??):159–160, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300546.

Mousavi:2017:MFR

[MS17] Seyed Nourollah Mousavi and Helle Sørensen. Multinomial functional regression with wavelets and LASSO penalization. Econometrics and Statistics, 1(??):150–166, January 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300211.

Meligkotsidou:2017:BAU

[MTV17] Loukia Meligkotsidou, Elias Tzavalis, and Ioannis Vrontos. On Bayesian analysis and unit root testing for autoregressive models in the presence of multiple structural breaks. Econometrics and Statistics, 4(??):70–90, October 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306217300369.

Muller:2018:HNE

[Mul18] Christophe Muller. Heterogeneity and nonconstant effect in two-stage quantile regression. Econometrics and Statistics, 8 (??):3–12, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300655.

Mutschler:2018:HOS

[Mut18] Willi Mutschler. Higher-order statistics for DSGE models. Econometrics and Statistics, 6(??):44–56, April 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300077.

McElroy:2020:MLP

[MW20] Tucker S. McElroy and Marc Wildi. The multivariate linear prediction problem: Model-based and direct filtering solutions. Econometrics and Statistics, 14(??):112–130, April 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300034.

Mukhopadhyay:2023:PRS

[MW23] Subhadeep Mukhopadhyay and Kaijun Wang. On the problem of relevance in statistical inference. Econometrics and Statistics, 25(??):93–109, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001283.

Manisera:2022:MMO

[MZ22] Marica Manisera and Paola Zuccolotto. A mixture model for ordinal variables measured on semantic differential scales. Econometrics and Statistics, 22(??):98–123, April 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000782.

Niessl:2023:SIS

[NABM23] Alexandra Nießl, Arthur Allignol, Jan Beyersmann, and Carina Mueller. Statistical inference for state occupation and transition probabilities in non-Markov multi-state models subject to both random left-truncation and right-censoring. Econometrics and Statistics, 25(??):110–124, January 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001155.

Norkute:2019:FAM

[NW19] Milda Norkute and Joakim Westerlund. The factor analytical method for interactive effects dynamic panel models with moving average errors. Econometrics and Statistics, 11(??):83–104, July 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300716.

Orsal:2017:MAC

[ÖA17] Deniz Dilan Karaman Örsal and Antonia Arsova. Meta-analytic cointegrating rank tests for dependent panels. Econometrics and Statistics, 2(??):61–72, April 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306216300028.

Omerovic:2022:MMR

[OFG22] Sanela Omerovic, Herwig Friedl, and Bettina Grün. Modelling multiple regimes in economic growth by mixtures of generalised nonlinear models. Econometrics and Statistics, 22 (??):124–135, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000307.

Ombao:2024:SD

[OP24] Hernando Ombao and Marco Pinto. Spectral dependence. Econometrics and Statistics, 32(??):122–159, October 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306222001101.

Otto-Sobotka:2019:ASM

[OSSRK19] Fabian Otto-Sobotka, Nicola Salvati, Maria Giovanna Ranalli, and Thomas Kneib. Adaptive semiparametric M-quantile regression. Econometrics and Statistics, 11(??):116–129, July 2019. CO-

DEN ???? ISSN 2452-3062. URL https://www.sciencedirect.
com/science/article/pii/S245230621930019X.

Paolella:2017:ASP

[Pao17] Marc S. Paolella. Asymmetric stable Paretian distribution testing. Econometrics and Statistics, 1(??):19–39, January 2017. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306216300247.

Pigini:2022:CIB

[PB22] Claudia Pigini and Francesco Bartolucci. Conditional inference for binary panel data models with predetermined covariates. Econometrics and Statistics, 23(??):83–104, July 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000101.

Platoni:2020:HST

[PBMS20] Silvia Platoni, Laura Barbieri, Daniele Moro, and Paolo Sckokai. Heteroscedastic stratified two-way EC models of single equations and SUR systems. Econometrics and Statistics, 15 (??):46–66, July 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621930022X.

Phillip:2020:GBS

[PCP20] Andrew Phillip, Jennifer Chan, and Shelton Peiris. On generalized bivariate Student-t Gegenbauer long memory stochastic volatility models with leverage: Bayesian forecasting of cryptocurrencies with a focus on Bitcoin. Econometrics and Statistics, 16(??):69–90, October 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300753.

Peng:2024:SAN

[PL24] Rong Peng and Zudi Lu. Semiparametric averaging of nonlinear marginal logistic regressions and forecasting for time series classification. Econometrics and Statistics, 31 (??):19–37, July 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001325.

Pollock:2018:SPL

[Pol18] D. S. G. Pollock. Stochastic processes of limited frequency and the effects of oversampling. Econometrics and Statistics, 7(??):18–29, July 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300011.

Proietti:2023:SHF

[PP23] Tommaso Proietti and Diego J. Pedregal. Seasonality in high frequency time series. Econometrics and Statistics, 27 (??):62–82, July 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000090.

Panayi:2018:DMM

[PPDZ18] Efstathios Panayi, Gareth W. Peters, Jon Danielsson, and Jean-Pierre Zigrand. Designating market maker behaviour in limit order book markets. Econometrics and Statistics, 5 (??):20–44, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300193.

Paindaveine:2017:PTE

[PRV17] Davy Paindaveine, Rondrotiana Joséa Rasoafaraniaina, and Thomas Verdebout. Preliminary test estimation for multisample principal components. Econometrics and Statistics, 2 (??):106–116, April 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300060.

Pesaran:2019:BAL

[PS19a] M. Hashem Pesaran and Ron P. Smith. A Bayesian analysis of linear regression models with highly collinear regressors. Econometrics and Statistics, 11(??):1–21, July 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300728.

Poloni:2019:CFR

[PS19b] Federico Poloni and Giacomo Sbrana. Closed-form results for vector moving average models with a univariate estimation approach. Econometrics and Statistics, 10(??):27–52, April 2019. CODEN

???? ISSN 2452-3062. URL https://www.sciencedirect.com/
science/article/pii/S2452306218300327.

Pesaran:2023:APT

[PS23] M. Hashem Pesaran and Ron P. Smith. Arbitrage pricing theory, the stochastic discount factor and estimation of risk premia from portfolios. Econometrics and Statistics, 26 (??):17–30, April 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001362.

Psaradakis:2024:MSM

[PS24] Zacharias Psaradakis and Martin Sola. Markov-switching models with state-dependent time-varying transition probabilities.

Econometrics and Statistics, 29(??):49–63, January 2024. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306221000575.

Psaradakis:2017:DTN

[PV17] Zacharias Psaradakis and Marián Vávra. A distance test of normality for a wide class of stationary processes. Econometrics and Statistics, 2(??):50–60, April 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306216300296.

Petersen:2022:MPD

[PZK22] Alexander Petersen, Chao Zhang, and Piotr Kokoszka. Modeling probability density functions as data objects. Econometrics and Statistics, 21(??):159–178, January 2022. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S245230622100054X.

Quessy:2019:CCA

[QD19] Jean-François Quessy and Martin Durocher. The class of copulas arising from squared distributions: Properties and inference. Econometrics and Statistics, 12(??):148–166, October 2019. CO-

DEN ???? ISSN 2452-3062. URL https://www.sciencedirect.
com/science/article/pii/S2452306219300103.

Riani:2024:ICO

[RAC^{+}24] Marco Riani, Anthony Curtis Atkinson, Aldo Corbellini, Alessio Farcomeni, and Fabrizio Laurini. Information criteria for outlier detection avoiding arbitrary significance levels. Econometrics and Statistics, 29(??):189–205, January 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306222000107.

Rodriguez-Caballero:2022:ECG

[RC22] Carlos Vladimir Rodríguez-Caballero. Energy consumption and GDP: a panel data analysis with multi-level crosssectional dependence. Econometrics and Statistics, 23(??): 128–146, July 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000083.

Rich:2020:SFC

[Ric20] Jeppe Rich. A spline function class suitable for demand models. Econometrics and Statistics, 14(??):24–37, April 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306218300078.

Ronchetti:2020:ARI

[Ron20] Elvezio Ronchetti. Accurate and robust inference. Econometrics and Statistics, 14(??):74–88, April 2020. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306220300022.

Rousseeuw:2019:RMT

[RPRH19] Peter Rousseeuw, Domenico Perrotta, Marco Riani, and Mia Hubert. Robust monitoring of time series with application to fraud detection. Econometrics and Statistics, 9(??): 108–121, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300303.

Rombouts:2020:VSP

[RSV20] Jeroen V. K. Rombouts, Lars Stentoft, and Francesco Violante. Variance swap payoffs, risk premia and extreme market conditions. Econometrics and Statistics, 13(??):106– 124, January 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300279.

Stoehr:2021:DCC

[SAK21] Christina Stoehr, John A. D. Aston, and Claudia Kirch. Detecting changes in the covariance structure of functional time series with application to fMRI data. Econometrics and Statistics, 18(??):44–62, April 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300460.

Sahin:2022:VCM

[SC22] Özge Sahin and Claudia Czado. Vine copula mixture models and clustering for non-Gaussian data. Econometrics and Statistics, 22(??):136–158, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001052.

Segers:2017:NAM

[SFdB17] Rene Segers, Philip Hans Franses, and Bert de Bruijn. A novel approach to measuring consumer confidence. Econometrics and Statistics, 4(??):121–129, October 2017. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306216300338.

Sakaria:2017:EBI

[SG17] D. K. Sakaria and J. E. Griffin. On efficient Bayesian inference for models with stochastic volatility. Econometrics and Statistics, 3(??):23–33, July 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300090.

Shang:2017:FTS

[Sha17] Han Lin Shang. Functional time series forecasting with dynamic updating: An application to intraday particulate matter concentration. Econometrics and Statistics, 1(??): 184–200, January 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306216300235.

Skripnikov:2019:JEM

[SM19] A. Skripnikov and G. Michailidis. Joint estimation of multiple network Granger causal models. Econometrics and Statistics, 10(??):120–133, April 2019. CODEN ???? ISSN

2452-3062. URL https://www.sciencedirect.com/science/
article/pii/S2452306218300509.

Sampaio:2020:SRG

[SM20] Jhames M. Sampaio and Pedro A. Morettin. Stable randomized generalized autoregressive conditional heteroskedastic models. Econometrics and Statistics, 15(??):67–83, July 2020. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306218300947.

Smith:2023:ICO

[Smi23] Michael Stanley Smith. Implicit copulas: an overview. Econometrics and Statistics, 28(??):81–104, October 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001465.

Seri:2021:MCV

[SMSC21] Raffaello Seri, Mario Martinoli, Davide Secchi, and Samuele Centorrino. Model calibration and validation via confidence sets. Econometrics and Statistics, 20(??):62–86, October 2021. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306220300162.

Shirota:2017:CRS

[SOLP17] Shinichiro Shirota, Yasuhiro Omori, Hedibert. F. Lopes, and Haixiang Piao. Cholesky realized stochastic volatility model. Econometrics and Statistics, 3(??):34–59, July 2017. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300181.

Savinov:2023:RTT

[SS23] Evgeniy Savinov and Victoria Shamraeva. On a Rosenblatt-type transformation of multivariate copulas. Econometrics and Statistics, 25(??):39–48, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001313.

Song:2023:BAA

[SSL23] Zefang Song, Xinyuan Song, and Yuan Li. Bayesian analysis of ARCH-M model with a dynamic latent variable. Econometrics and Statistics, 28(??):47–62, October 2023. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221001167.

Sass:2024:RRP

[ST24a] Jörn Sass and Anna-Katharina Thös. Risk reduction and portfolio optimization using clustering methods. Econometrics and Statistics, 32(??):1–16, October 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001416.

Silvennoinen:2024:CAN

[ST24b] Annastiina Silvennoinen and Timo Teräsvirta. Consistency and asymptotic normality of maximum likelihood estimators of a multiplicative time-varying smooth transition correlation GARCH model. Econometrics and Statistics, 32(??): 57–72, October 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000848.

Simone:2020:SHR

[STI20] Rosaria Simone, Gerhard Tutz, and Maria Iannario. Subjective heterogeneity in response attitude for multivariate ordinal outcomes. Econometrics and Statistics, 14(??):145–158, April 2020. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306219300243.

Schmid:2018:DMD

[STW18] Matthias Schmid, Gerhard Tutz, and Thomas Welchowski. Discrimination measures for discrete time-to-event predictions. Econometrics and Statistics, 7(??):153–164, July 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306217300321.

Sujica:2018:CGE

[SV18] Aleksandar Sujica and Ingrid Van Keilegom. The copulagraphic estimator in censored nonparametric location-scale regression models. Econometrics and Statistics, 7(??):89–114, July 2018. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306217300631.

Sutton:2019:MIR

[SVG19] Maxwell Sutton, Andrey L. Vasnev, and Richard Gerlach. Mixed interval realized variance: a robust estimator of stock price volatility. Econometrics and Statistics, 11(??):43–62, July 2019. CODEN ???? ISSN 2452-3062. URL https://www.

sciencedirect.com/science/article/pii/S2452306218300285.

Schatz:2022:APP

[SWS22] Michael Schatz, Spencer Wheatley, and Didier Sornette. The ARMA point process and its estimation. Econometrics and Statistics, 24(??):164–182, October 2022. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001349.

Sun:2019:ESV

[SZH19] Yanqing Sun, Yuanqing Zhang, and Jianhua Z. Huang. Estimation of a semiparametric varying-coefficient mixed regressive spatial autoregressive model. Econometrics and Statistics, 9(??):140–155, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300424.

Tang:2024:MST

[Tan24] Cheng Yong Tang. A model specification test for semiparametric nonignorable missing data modeling. Econometrics and Statistics, 30(??):124–132, April 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230622100099X.

Tutz:2017:SLD

[TB17] G. Tutz and M. Berger. Separating location and dispersion in ordinal regression models. Econometrics and Statistics, 2 (??):131–148, April 2017. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621630003X.

Tutz:2020:EEV

[TB20] Gerhard Tutz and Moritz Berger. The effect of explanatory variables on income: a tool that allows a closer look at the differences in income. Econometrics and Statistics, 16 (??):28–41, October 2020. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218301035.

Tong:2018:ECT

[THS18] Xiaojun Tong, Zhuoqiong Chong He, and Dongchu Sun. Estimating Chinese Treasury yield curves with Bayesian smoothing splines. Econometrics and Statistics, 8(??):94–124, October 2018. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306217300898.

Takahashi:2024:FDV

[TWO24] Makoto Takahashi, Toshiaki Watanabe, and Yasuhiro Omori. Forecasting daily volatility of stock price index using daily returns and realized volatility. Econometrics and Statistics, 32 (??):34–56, October 2024. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000964.

Villani:2024:SSM

[VQKS24] Mattias Villani, Matias Quiroz, Robert Kohn, and Robert Salomone. Spectral subsampling MCMC for stationary multivariate time series with applications to vector ARTFIMA processes. Econometrics and Statistics, 32(??):98–121, October 2024. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S245230622200106X.

Vallejos:2017:IUH

[VS17] Catalina A. Vallejos and Mark F. J. Steel. Incorporating unobserved heterogeneity in Weibull survival models: A Bayesian approach. Econometrics and Statistics, 3(??):73–88, July 2017. CO-

DEN ???? ISSN 2452-3062. URL http://www.sciencedirect.
com/science/article/pii/S2452306217300072.

Voges:2021:CFC

[VS21] Michelle Voges and Philipp Sibbertsen. Cyclical fractional cointegration. Econometrics and Statistics, 19(??):114–129, July 2021. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306220300502.

Wee:2022:LIM

[WCD22] Damien C. H. Wee, Feng Chen, and William T. M. Dunsmuir. Likelihood inference for Markov switching GARCH(1,1) models using sequential Monte Carlo. Econometrics and Statistics, 21(??):50–68, January 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300368.

Wang:2018:CQR

[WCW18] Meng Wang, Zhao Chen, and Christina Dan Wang. Composite quantile regression for GARCH models using high-frequency data. Econometrics and Statistics, 7(??):115–133, July 2018. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306216300284.

Wagner:2023:FAB

[WFSJ23] Helga Wagner, Sylvia Frühwirth-Schnatter, and Liana Jacobi. Factor-augmented Bayesian treatment effects models for panel outcomes. Econometrics and Statistics, 28(??): 63–80, October 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306222000430.

Wenger:2021:FBC

[WL21] Kai Wenger and Christian Leschinski. Fixed-bandwidth CUSUM tests under long memory. Econometrics and Statistics, 20 (??):46–61, October 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306219300474.

Wu:2018:SEU

[WS18] Ximing Wu and Robin Sickles. Semiparametric estimation under shape constraints. Econometrics and Statistics, 6 (??):74–89, April 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300436.

Werge:2022:AAR

[WW22] Nicklas Werge and Olivier Wintenberger. AdaVol: an adaptive recursive volatility prediction method. Econometrics and Statistics, 23(??):19–35, July 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000113.

Xue:2022:MLE

[XY22] Jiacheng Xue and Weixin Yao. Machine learning embedded semiparametric mixtures of regressions with covariatevarying mixing proportions. Econometrics and Statistics, 22 (??):159–171, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001453.

Yoshida:2018:SMM

[Yos18] Takuma Yoshida. Semiparametric method for model structure discovery in additive regression models. Econometrics and Statistics, 5(??):124–136, January 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S245230621730014X.

Sin:2021:UHN

[ySL21] C. Y. (Chor yiu) Sin and Cheng-Few Lee. Using heteroscedasticitynon-consistent or heteroscedasticity-consistent variances in linear regression. Econometrics and Statistics, 18(??):117–142, April 2021. CODEN ???? ISSN 2452-3062. URL http://www.

sciencedirect.com/science/article/pii/S2452306220300848.

Yi:2023:MDP

[YW23] Yanqing Yi and Xikui Wang. A Markov decision process for response adaptive designs. Econometrics and Statistics, 25 (??):125–133, January 2023. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221001301.

Zhuang:2022:BNM

[ZDY22] Haoxin Zhuang, Liqun Diao, and Grace Y. Yi. A Bayesian nonparametric mixture model for grouping dependence structures and selecting copula functions. Econometrics and Statistics, 22(??):172–189, April 2022. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306221000526.

Zhong:2024:ESM

[ZHO24] Peng Zhong, Raphaël Huser, and Thomas Opitz. Exact simulation of max-infinitely divisible processes. Econometrics and Statistics, 30(??):96–109, April 2024. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306222000156.

Zhao:2018:TDM

[ZMG18] Yixing Zhao, Rogemar Mamon, and Huan Gao. A twodecrement model for the valuation and risk measurement of a guaranteed annuity option. Econometrics and Statistics, 8 (??):231–249, October 2018. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306218300340.

Zou:2021:BSU

[ZP21] Nan Zou and Dimitris N. Politis. Bootstrap seasonal unit root test under periodic variation. Econometrics and Statistics, 19(??):1–21, July 2021. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306220300174.

Zhang:2024:STA

[ZPY^{+}24] Tingting Zhang, Minh Pham, Guofen Yan, Yaotian Wang, Sara Medina-DeVilliers, and James A. Coan. Spatial-temporal analysis of multi-subject functional magnetic resonance imaging data. Econometrics and Statistics, 31(??):117–129, July 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306221000289.

Zhou:2024:EHD

[ZY24] Yunpeng Zhou and Kam Chuen Yuen. Estimation in the high dimensional additive hazard model with l_{0} type of penalty. Econometrics and Statistics, 32(??):88–97, October 2024. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306222000831.

Zhang:2022:NGU

[ZYL22] Yixiao Zhang, Cindy L. Yu, and Haitao Li. Nowcasting GDP using dynamic factor model with unknown number of factors and stochastic volatility: a Bayesian approach. Econometrics and Statistics, 24(??):75–93, October 2022. CODEN ???? ISSN

2452-3062. URL http://www.sciencedirect.com/science/
article/pii/S2452306221001039.

Zhang:2019:ETI

[ZZ19] Yonghui Zhang and Qiankun Zhou. Estimation for timeinvariant effects in dynamic panel data models with application to income dynamics. Econometrics and Statistics, 9 (??):62–77, January 2019. CODEN ???? ISSN 2452-3062.

URL http://www.sciencedirect.com/science/article/pii/
S2452306217300904.

Zhang:2020:SSI

[ZZZ20] Liang Zhang, Tianming Zhu, and Jin-Ting Zhang. A simple scale-invariant two-sample test for high-dimensional data. Econometrics and Statistics, 14(??):131–144, April 2020. CODEN

???? ISSN 2452-3062. URL http://www.sciencedirect.com/
science/article/pii/S2452306220300010.